A Selected Bibliography of Publications by, and about, John Clarke Slater

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.73

Title word cross-reference

#1 [Sla62].

$10 [Sla56b]. 14.5 × 22 [Sla56b]. $5.95 [Sla68b]. ^{+} [Pra52, SWW69]. ^{+2} [WWS70]. ^{−} [Sla59c]. _{2} [Sla41, Sla87]. _{3} [Sla31c]. _{3}^{−} [Sla59c]. _{4} [Sla41, Sla87]. α [Sla73a]. α [Con77, Sla70b, Sla72a, SJ72, Sla74b, Sla75a]. d [RFST85]. sech^{2} [HMMT04].

11.50 [OT80]. 1937 [Ano37]. 1949 [Sla49e]. 1956 [Sla56f]. 1966 [Wel67]. 1967 [Ano67, Löw67]. 1971 [Sla71b]. 1973 [DP74, Sla73b]. 1976 [Ame82, LW77, Mor82]. 1989 [Mol95].

3-cm [BGH^{+}46]. 387pp [Sla56b].

4 [Cho74]. 44.85 [T.75].

50th [PCR73].

60s [Con10].

7 [KMV77, Wei70a].

Absorption [HS25, Sla25b, Sla27b, SS36]. Accelerator [MS52b, TS52, DKS52]. Accelerators [Sla48a, Sla52b]. Accuracy [Par09, RLER04]. Action [Sla27a]. Administrative [Sla62]. Again [Sla49d]. age [Van64]. Alarm [Rid84]. Alkali [Sla23, Sla24a, Sla26c, SS36]. Alkali-Halides [Sla23]. Alla [Sla49b]. Alloys [Sla37g, Sla40b, Sla56d, Sla56b]. Alma [CWH^{+}52]. Alternating [Sla26a]. ALTRAN [FP73]. America [Sla67c, Sla68g, Sop76, Sop80, Sop88]. American [MJW71, Smi81, Wil76, Van64, Rom91]. Americans [Jan08]. Amos [CWH^{+}52]. Analysis [LS70]. Analytic [Sla32a]. Anniversary [PCR73]. Anomalous [BU26]. anti [Sla70a]. anti-symmetric [Sla70a]. Antisymmetric [Fol62]. Application [Pau31a]. Applications [ERW75, MMSW97, PW35, Seg77b]. appraisal [RFST85]. Approach [SHR13, PAAC93]. Approximation [Gás54, Pra52, WWS70]. APW [Sla67d]. Aspects [CWH^{+}52]. Atom [Sla27c]. Atomic [CC61, Com56, HS63, Löw67, Sla30a, Sla32a, Sla51h, Sla60c, Sla63a, Sla64c, Sla68b, Sla71a, Meg61]. Atoms [CS18, Hof67, Löw66b, Sla24b, Sla26b, Sla27a, Sla28d, Sla53d, Sla55a, Sla63b, Sla68a, SMWW69, WWS70, PAAC93, Sla50b, Sla56d, Sla56b, Sla71d, Sla76b]. Attempt [Kon98]. Augmented [SS53, Sla53a, Sla64b, Sla67d, Woo62]. August [Mol95, Wei70a]. Autonomy [Hen03]. Average [Sla68a].

B [Seg77b]. balanced [PAAC93]. Baldwin [CWH^{+}52]. Balmer [Sla20]. Baltimore [Sla68b]. Band [Joh65, MJW71, Sla26a, Sla53b, Sla56c, Sla59a, Sla64b, Sla66, Sla68d, Sla53e, Sla68i]. Bands [AFR68, Sla34a, Sla52a, SKW62, Sla65a, Sla68c, Sla72e, Woo62, Sla67b, Sla68f, Sla68e]. Barium [Sla50a]. Barrier [Sla56a]. Based [WFS^{+}01, RFST85]. Basic [Sla76a]. basis [PAAC93, RLER04]. Becoming [Kai10]. beginner [Sla67d]. beginnings [Con10]. Behavior [SS52, UG25]. Bernard [CWH^{+}52]. Beta [Joh65]. Beta-Brass [Joh65]. Between [Amo97, CS18, Par09, Pau40, Sla67c, Sla68g, USOA13]. beyond [HM05]. bezüglich [UG25]. Bible [CWH^{+}52]. Bibliographies [SYLxx]. Big [Hil15]. Biochemistry [CWH^{+}52]. Biographical [Mor82]. Biography [Sla75d, Ehr75, Sop78]. Biology [LP64, CWH^{+}52]. Birth [Hil15]. bodies [Sla38a]. Bohr [BG24, BG24, FK85, Hen81, Kle70, Kon83, Kra09, Was81]. Bomb [Gol92]. Bond [Pau31a, Pau31b]. bonding [Sla56c]. Bonds [Sla31b, Dus36]. Book [Ano53, Bar56, Bli69, CC61, Chi64, Cho74, Coh56, CWH^{+}52, Dew63, Dus52, Ehr75, Ett35a, Fer48, Gar70, Hay51, Hof67, Kle53, Kon80a, Kon80b, Meg61, Mer69, Pag34, Phi74, Sla56d, Sla56b, Sla64a, Sla67d, Sla68b, Sla75b, Sop78, Tay69, Tol56, Wit52]. Bose [Kos86]. both [RFST85]. Boulder [SM73]. Bradford [Mol95]. Brass [Joh65]. Breadth [Kon98]. Breadths [HS25]. Brian [Sla68b]. Bridgman [Sla65d]. Brief [Löw85]. Brillouin [Sla51d]. Britain [Rid84]. British [CWH^{+}52]. Broglie [PCR73]. Brother [CWH^{+}52]. Build [Gol92]. Bush [Gol92].

C [Ano53, Ano67, Anoxx, Ano23, Bar56, Bli69, Bom01, CC61, Chi64, Cho74, Coh56, Con10, CWH^{+}52, Dew63, Dus52, Ehr75, Ett35a, Ett35b, Fer48, Gar70, Hay51, Hof67, Jon76, Kle53, Kon80a, Kon80b, Kon97, KMV77, Löw66a, Löw66b, Meg61, Mer69, Nes10, OT80, Pag34, Phi74, Sop78, T.75, Tay69, Tol56, Van76, Wit52]. C. [Sla64a]. Calculation [HCST73, Kon80a, Kon80b, KS54a, Seg77a, Seg77b, Sla79, Hat72, OT80]. Calculations [HS63, Sla64b, Sla65e, Sla67d, Sla76b, Sla67d]. Capturing [Moh13]. Casting [CWH^{+}52]. catalysis [SJ74]. Cavities [MS52a]. Cavity [MS52b, SS52]. cellular [Sla76b]. Centenary [FK85]. Center [MJW71, Sla60b]. Central [Sla28a]. Change [Mar15]. changed [Sla69b]. Changing [Par03]. Chapters [HBTW92]. Charge [WFS^{+}01]. Charles [Sla37f]. Chauncey [CWH^{+}52]. Chemical [Pau31a, Pau31b, Sla39, Sla51a, Sla70c]. Chemistry [DP74, Ehr75, LP64, MMSW97, Par03, Par09, PW35, Sch90, Sla49b, Sla56b, SHR13, SJ74, SM73]. Chemists [Hen03, Hen04]. Chimica [Sla49b]. Clarke [Ame82, And77, Anoxx, Bom01, Hoc90, Hor00, Löw67, Mor82, Sch90, Wei70a, Wei07, Wil76, Hod08]. Claude [CWH^{+}52]. Climate [Gol92]. Cluster [SJ72]. cm [BGH^{+}46, Sla56b]. Coding [Fis19]. Coefficients [Löw83]. Cohesion [Löw66a, Sla30b]. College [LW77]. Combination [Sla51a]. comes [Van64]. Comments [Löw66a]. Committee [Sla52c]. Commonwealth [CWH^{+}52]. Comparative [Wag83]. Comparison [Mag14, SWW69, Sla75a]. Compendium [GHW09]. Complex [Hil15, Sla26b, Sla29a]. Compressibility [Sla23, Sla24a, Sla26c]. Computation [ISS06, USOA13, ZD11]. Computational [MJW71, Par09, SHR13]. Computers [Par03]. computing [Don11]. Concepts [GHW09, Sla69c, Gar70]. Condon [Sla73b, BO71, VV91]. Conduction [Sla33a]. Conference [LW77, MJW71, Sla49e, Sla58a, SM73]. Configuration [Sla54]. configurations [Tud93]. Confocal [SS52]. Congress [DP74]. Connection [Pau40]. Connolly [SWS^{+}73]. Considerations [Sla37h]. Consistent [KS65, Phi74, Sla28d, Sla53c, Sla69a, SJ72, Sla72d, Sla74c, Sla51i, Sla71c, Sla78]. Constant [Sla40a]. Constants [Sla30a]. Contexts [Par00]. Conversion [Sla51a]. Cornell [Ano37]. Correction [Sla50a]. Correlation [KS65, Sla63b, Sla67a, Sla72d]. Crystal [HBTW92, Sla58a, SW70, SW71]. Crystals [Sla37a, Sla58b, Sla64c, Sla65a, Sla65b, SMWW69, Sla69a, Sla72e, Sla38b, Sla71c]. Cu [Pra52, SWW69]. Curie [Sla40b]. Curmudgeon [And11a]. Current [Sla58a, Sla67e]. Cusp [Moh13]. Cusps [RH09]. Cyclotron [Hil15].

D [CWH^{+}52, SWS^{+}73]. Damped [Sla37a]. December [Mor82]. Decision [Gol92, Kai10]. Degenerate [Sla25d]. della [Sla80]. densities [BT88]. Density [LS70, Con10, HMMT04, Kos86, RH09]. Department [Sla48b]. Dependence [Sla20]. Depression [Wei70b]. derivatives [RH09]. Design [LC10, Sla48a]. Determinantal [Sla51f]. Determinants [Amo97, Fol62, Löw83, Moh13, USOA13]. Determination [MS52b]. Determining [Sla25b]. Development [Gar70, Sch90, Sla69c, Was81, Sla73c]. Dialogue [Kle70]. Different [And11a]. Digital [Par03]. dimensional [HMMT04]. Dirac [ERW75, Kos86, Tal04]. Directed [Sla31a]. Discipline [Par03]. Discovery [Par00, PCR73]. Discussion [Har65, Sla63a, SWS^{+}73, Sla65c]. Disorder [Sla37h]. dispersion [Kon97]. DM [T.75]. Doublets [BU26]. Durch [Gás54, UG25]. Dutch [GU26]. Dynamic [CWH^{+}52]. Dynamical [Sla26b]. Dynamics [Sla25d, TS52, Don11, ZD11].

E. [Sla73b]. each [UG25]. Early [Par03]. Earnest [CWH^{+}52]. edition [Sla56b]. Editorial [Rom91]. Edward [BO71]. Effect [BU26, Sla40b, Sla51a]. Effective [WFS^{+}01]. Effects [KS65, Sla36b, Sla51g, Sla51c, PAAC93]. Efficient [USOA13]. Egyptian [CWH^{+}52]. Einstein [Kle70, Kos86]. einzelnen [UG25]. Electrodynamics [Sla38a]. Electromagnetic [SS52]. Electromagnetism [SF47a, SF69, SF73]. Electron [BU26, DKS52, Pau31b, PCR73, Sla33a, Sla37a, SSK53, Sla55a, SKW62, Tho26, GU26, Sla51b, Sla70a, Tud93, UG25, LW77]. Electronic [Chi64, Joh65, Par03, Seg77a, Seg77b, Sla34a, Sla34b, Sla37g, Sla53d, Sla53e, Sla53f, Sla54, Sla63b, Sla63c, Sla67d, Dew63, Sla57]. Electronic-Structure [Sla63c]. Electronics [Sla46a, Sla50c, Sla69d, Hay51]. Electrons [RS36, Sla26d, Sla49a, SWW69, UG26, Sla56d, Sla56b]. electrostatic [RLER04]. Elektrons [UG25]. Elements [Amo97, Dus36, USOA13]. Elwood [Sla37f]. Emigration [Rid84]. Employing [Fis19]. Enemy [CWH^{+}52]. Energies [Sla55a]. Energy [AFR68, Pra52, Sla31b, Sla34a, Sla37d, Sla52b, Sla52a, SKW62, Sla64b, Sla65a, Sla66, Sla67b, Sla68a, Sla68f, Sla68d, Sla68e, Sla68c, SW70, SW71, Sla72e, WP57, Woo62, Hat72, HCST73, HMMT04, Sla38b, Sla53e, Sla68i, SM73]. Energy-Band [Sla66]. Engineers [CWH^{+}52]. Equation [Sla31d, Sla51c, Sla46b]. Equations [KS65, Pra52]. Ernest [Hil15]. Ersetzung [UG25]. Establishment [Cob72]. Ettore [EAW15]. Evaluation [Amo97, DWZ22, FP73]. Evans [CWH^{+}52]. Evidence [Kon83]. Exact [FP73]. Example [SSK53]. Excel [PM12]. Exchange [KS65, Sla68c, SWW69, SW70, SW71, Sla71a, Sla72d, WWS70, HM05, Sla70b, Sla72c]. Exchange-Correlation [Sla72d]. Excited [Sla38b]. Executive [Sla52c]. Expanded [Sla56f, RH09]. Expansion [Löw83]. Experimental [CWH^{+}52, BG24]. experimentellen [BG24]. Experiments [GHW09]. Explanation [BU26, Hen03, Hen04]. Extension [Gou30].

F [CWH^{+}52, Ett35a]. F. [Ett35b, Sla75b]. Fast [ISS06]. Fe [WP57]. February [KMV77, Mol95, Wei70a]. Fellow [Sla56b]. Fermi [Kos86, SK35, Sla67b]. Ferroelectricity [Wel67, Sla67f, Wel67]. Ferromagnetism [Sla36a, Sla36b, Sla37d, Sla53b, SSK53, Sla53g]. Feynman [AM06, Sla72a]. FHF [Sla59c]. Field [MS52b, MS52a, Phi74, Sla28d, Sla53c, Sla69a, SJ72, Sla72d, Sla74c, Sla78, AM06, Sla71c]. Fields [Sla28a]. Fifty [Ehr75, PCR73]. figs [Sla56b]. Film [Gar70]. Fireside [CWH^{+}52]. First [DP74, Kle70, PCR73, WWS70]. First-Transition-Row [WWS70]. Fishery [CWH^{+}52]. Fisica [Sla49b, SF51]. Fits [PM12]. Fitting [PM12]. Florida [Löw67]. Fock [Wag83, Gás54, LS70, Pra52, Sla51c, Sla51e, Sla63b, Wag83, WP57]. Fockschen [Gás54]. Following [SWS^{+}73, Har65]. force [UG25]. Forces [SK31]. Forderung [UG25]. Formulas [KU94, Sla28a]. Found [Pra52, WP57]. fractional [Sla71d]. France [DP74]. Frank [Ett35a, Ett35b, Fer48, Pag34]. Free [SKW62]. Freshwater [CWH^{+}52]. Function [Joh65, Sla65b, Sla66, Gás54, Hat72, HCST73]. functional [Con10]. Functions [Fol62, KS54b, KS54c, Pra52, Sla32a, Sla37e, SHR13, WP57, Kos86, PAAC93, RH09, Sla51f, Sla70a]. Future [Sla73a].

G [CWH^{+}52]. Gallium [SKW62]. Gases [SK31]. Gaussian [BT88, Mag14, PAAC93, RFST85, RLER04, SHR13]. Gaussian-Type [Mag14]. Gaussians [PM12]. Geminal [Moh13]. General [KU94, Wel67]. Generalized [Sla53c, VV91]. Generation [Sla75c]. Genesis [Kon83]. Genius [Ree99]. German [BKS24b, BG24, Gás54, UG25]. Given [Sla68a]. graphics [DWZ22]. Great [Wei70b]. Green [Joh65]. Green’s [Sla66]. Green’s-Function [Sla66]. Ground [PAAC93, Don11, ZD11]. Group [Sla53h, Sla53i]. Groups [Sla31c, Sla65b]. Grüneisen [Sla40a]. Guide [Sla67d].

H [Ett35a, Ett35b, Fer48, Pag34, Sla71b, SWS^{+}73, AM06]. Halides [Sla23, Sla24a, Sla26c, SS36]. Hamiltonian [AM06, USOA13]. Harold [CWH^{+}52]. Harris [CWH^{+}52]. Hartree [Gás54, WP57, Gás54, LS70, Pra52, Sla30c, Sla51c, Sla51e, Sla63b, Wag83]. Heights [MJW71]. Heisenberg [Sla72c]. Heitler [Sla65c]. Held [DP74, LW77, Löw67, MJW71]. Helium [Sla25a, Sla27c, Sla28c]. Hellmann [Sla72a]. High [Sla52b, Sla65d, RLER04]. High-Pressure [Sla65d]. high-quality [RLER04]. Hill [T.75]. Historic [Löw85]. History [GHW09, Hen81, HBTW92, Sla48b, Sla75b, WP85, Sla74b, Wei70a]. honor [Löw67]. Hopkins [Sla68b]. Houston [Gou30]. Howard [CWH^{+}52]. Huet [CWH^{+}52]. Hume [Sla56b, Sla56d]. Hume-Rothery [Sla56b, Sla56d]. Hund [Sla75b]. Hybridisation [Par00]. Hydrogen [CS18, Sla25a, PAAC93, Sla20]. hydrogenic [PAAC93]. Hyper [LS70]. Hyper-Hartree [LS70]. Hyperbola [Par03]. Hypothese [UG25]. hypothesis [UG25].

IBM [MJW71, MJW71]. Identity [Par03]. II [Pau31b, SS53, Sla36b, Sla37c, Sla53f, CC61]. III [Sla54]. Impedance [MMS49]. Imperatives [Par09]. Important [CDFS53a, CDFS53b, CDFS54]. improvements [Sla74a]. Impurity [KS54a, KS54b, KS54c]. Including [KS65]. Incompressible [Sla40a]. Industrial [Hil15]. Inelastic [RS36]. information [RFST85]. Injurious [CWH^{+}52]. inner [UG25]. inneren [UG25]. Inorganic [LC10]. Insects [CWH^{+}52]. Institute [Sla56f]. insulating [Sla38b]. Insulators [Mer69, Sla67g]. integral [Sla72c]. Integrals [ISS06, DWZ22, FP73, Tal04]. intense [AM06]. Intensities [Sla26a, Sla20]. Interaction [Sla58b, Sla54]. Interactions [CS18]. interatomic [Sla59c]. Interdisciplinary [Sla76a]. Interelectron [Moh13]. Internal [Sla51a]. International [DP74, LW77, Löw67, Sla49e, Sla58a]. Interpretation [Sla25a]. interview [Wei70a]. Introducción [SF51]. Introduction [Ett35b, Fis19, PW35, SF33, Sla39, Sla67f, Sla70c, Sla71b, Sla73b, Sla49b, SF51, Ett35a, Pag34]. Introduzione [Sla49b]. Inventing [Gol92]. Ion [SWW69, WWS70]. Ions [WWS70, Sla59c]. Iron [Sla40b, Woo62]. Irving [Ano67]. Island [Löw67]. Italian [Sla49b, Sla80]. Ithaca [Ano37]. Itinerant [LW77]. Itinerant-electron [LW77]. IUPAP [Sla52c]. IV [T.75].

J [CWH^{+}52, Ett35a, Ett35b, Fer48, Jon76, MJW71, Pag34, Sla71b]. J. [Kon97, SWS^{+}73]. January [Löw67, Sla71b, Sla73b]. jedes [UG25]. John [Anoxx, Bar56, Bli69, CC61, Chi64, Cho74, CWH^{+}52, Dew63, Dus52, Ehr75, Gar70, Hay51, Hod08, Hof67, Hor00, Kle53, Kon80a, Löw66a, Löw67, OT80, T.75, Wei07, Wil76, Wit52, Ame82, And11b, And77, Ano53, Ano67, Ano23, Bom01, Coh56, Con10, Hoc90, Kon80b, KMV77, Löw66b, Meg61, Mer69, Mor82, Nes10, Phi74, Sch90, Smi81, Sop78, Tay69, Van76, Van77, Wei70a]. Johns [Sla68b]. Johnson [SWS^{+}73]. joint [MJW71]. Journal [Rom91]. Judd [Sla68b]. July [Ano37, DP74, Mor82, Sla56f].

K. [SWS^{+}73]. Kahn [Kos86]. Karl [CWH^{+}52]. KH [Sla41, Sla87]. kinetic [HMMT04]. kinetically [PAAC93]. Kirkwood [CS18]. Kittel [Sla64a]. Koster [KU94]. Kramers [BG24, Hen81, Kon83, Kra09, Was81, BG24].

L [CWH^{+}52]. Laboratories [Sla76a, Wel67]. Lagler [CWH^{+}52]. Langmuir [Ano67]. language [FP73]. lattice [Hat72, HCST73, Sla53f]. Lattices [Sla49a]. Laureate [PCR73]. Lawrence [Hil15]. Layer [Moh13]. LCAO [PAAC93, SK54]. Lead [BGH^{+}46, Sla56a]. Leake [CWH^{+}52]. Least [PM12]. Least-Squares [PM12]. Lecturer [Sla56b]. Length [BGH^{+}46]. Levels [KS54b, KS54c, Pra52, Sla31b, Sla37d, WP57, Sla38b]. Library [Sla56b]. Light [Sla28b]. Line [HS25, Kon98, Sla25b]. Linear [DKS52, MS52b, Sla48a, Sla52b, TS52]. Lines [Sla26a, Sla20]. London [Sla65c]. Lorentz [Sla50a]. Loucks [Sla67d]. Louis [PCR73]. Low [Sla49e]. Löwdin [CGLÖ76, Hof67]. Lowest [Sla37d]. Lucky [Sla75c].

M [CWH^{+}52]. M.I.T. [DKS52, Sla48b]. Magnetic [BU26, Sla51c, AM06, Sla51f, And11b]. Magnetism [LW77, Sla68d]. magnetization [Hat72, HCST73]. Major [Gar70]. Majorana [EAW15]. Manageability [Par09]. manipulation [FP73]. many [Sla70a]. many-electron [Sla70a]. Maple [PM12]. Marcel [CWH^{+}52]. Massachusetts [Sla56f]. materia [Sla80]. Materials [LC10, Sla51g, Sla56e, Sla56f, Sla60b, Sla76a]. Mathematical [EAW15]. Mathematicians [Rid84]. Mathematics [CWH^{+}52]. Matlab [Fis19]. Matrix [Amo97, Jan08, LS70, USOA13]. Matter [Sla31e, Sla51j, Sla53j, Sla68h, Sla77, Sla80, Ano53, CWH^{+}52, Dus52, Kle53, Tay69, Wit52, Bli69]. May [MJW71]. Maze [HBTW92]. McGraw [T.75]. meaning [Sla29b]. Means [BU26]. Measurement [Sla26c]. Measurements [MS52a]. Mechanics [Cob72, Hen81, Hen03, Jan08, Pau31a, PW35, PCR73, Sla28a, Sla28b, SF47b, SF83, vdW67, Sla29b, Sla73c, Gar70, Fer48]. Medical [CWH^{+}52]. Medicine [CWH^{+}52]. Megacycles [MMS49]. Memo [Sla62]. Memoir [Mor82]. Memorable [Rom91]. Memorial [KMV77]. Mendenhall [Sla37f]. Menton [DP74]. Metallic [Ano37, Sla33a]. Metallurgical [Sla56b]. Metallurgy [CWH^{+}52]. Metals [Löw66a, Sla30b, Sla34a, Sla34b, SK35, Sla40a, Sla67g, Sla49c, Sla56d, Sla56c, Sla56b, Mer69]. Method [CS18, Con77, CWH^{+}52, Joh65, LS70, SS53, Sla30c, SK35, Sla51e, Sla53a, Sla53c, SK54, Sla63b, Sla64b, Sla66, Sla67d, Sla72a, SJ72, WP57, Woo62, Sla53e, Sla71c, Sla73a, Sla74b, Sla74a, SC76, Wag83]. Methods [CGLÖ76, MJW71, MMSW97, Seg77a, Seg77b, Sla25b, Sla65c, Don11, Sla72b, Sla75a, Sla76b]. Michigan [Wel67]. Microsoft [PM12]. Microwave [SS52, Sla42, Sla46a, Sla50c, Sla59b, Sla60a, Sla69d, Hay51]. Military [Hil15]. Mitchell [CWH^{+}52]. Mn [WWS70]. Model [Kon98, Sla26b, AM06, ERW75]. Modern [BO71, Sla55b, Coh56, Bar56, Tol56]. Moldenke [CWH^{+}52]. Molecular [Ehr75, Hen03, Kon80a, Kon80b, LP64, Löw67, Sla31b, Sla32b, Sla33b, Sla65c, Sla65e, Sla67e, Sla71a, Sla75d, Sla79, Sop78, OT80, Sla51i, Sla53h, Sla53i, SC76]. Molecules [Chi64, Cho74, Hof67, Löw66b, Mer69, Pau31a, Phi74, Sla31a, Sla53d, Sla55a, Sla63c, Sla65a, Sla67g, SJ72, Sla74c, Sla78, Dew63, ERW75, PAAC93, Sla50b, Sla72b, Sla75a, Sla76b, T.75, Chi64, Dew63]. moments [Kai10]. momentum [RFST85]. Monday [Sla71b]. Monovalent [Löw66a, Sla30b]. Motion [Tho26]. motivated [AM06]. Motors [Wel67]. Mott [And11b]. muffin [Sla74a]. muffin-tin [Sla74a]. Mulliken [Sla64d, LP64]. Multiplet [Gou30]. Multiplicities [Sla68a]. My [CWH^{+}52, CWH^{+}52].

N [CWH^{+}52, Ett35a, Ett35b, Fer48, Pag34]. Nachprüfung [BG24]. Name [Mar15]. narrative [Com56]. Nature [Pau31a, Pau31b, PCR73, Sla25c, Sla37b, Sla37c]. Nevill [And11b]. Newburyport [Sla64d]. NG [PM12]. Nickel [Sla36a, Sla36b, Sla40b]. Niels [FK85, Was81]. Nobel [PCR73]. Non [Amo97, USOA13]. Non-Orthogonal [Amo97, USOA13]. Nonintegral [SMWW69]. nonlocal [HM05]. Normal [MMS49, Sla28c]. Note [Sla30c, Sla31c, Sla32b, Sla40b, Sla40a, Sla51h, Sla51d, Sla59c, Sla70a, Don11]. Notes [And11a]. Nuclear [WFS^{+}01]. Numbers [SMWW69, Sla71d]. numerical [DWZ22]. NY [Ano37].

O [Sla31c]. Obituary [Jon76, Van76, Wil76]. Obtained [Pau31a]. Occupation [SMWW69, Sla71d]. oeuvre [And77]. Old [CWH^{+}52, Kon98]. Olmsted [CWH^{+}52]. Olov [Hof67, CGLÖ76]. One [Pau31b, Sla55a, HMMT04, BG24]. one-dimensional [HMMT04]. One-Electron [Pau31b, Sla55a]. Opinion [Gol92]. Opportunity [Rid84]. Optical [Sla25e, SS36, Kon97]. Oral [Wei70a]. Orbital [Sla65c]. Orbitals [LP64, Mag14, PM12, Sla51h, Sla79, RFST85, RH09, Sla51i, SC76, Tal04, Kon80a, OT80, Kon80b]. Order [Sla37h]. Ordered [Joh65]. Origins [Par00]. Orthogonal [Amo97, Sla51h, USOA13]. Oscillator [Kon98]. Oscillators [Hen81]. Overlap [Tal04]. overlapping [SC76]. overlapping-sphere [SC76]. Oxford [LW77, Sla56b].

P. [Sla65d]. Paper [Löw66a, SWS^{+}73, Was81, Har65]. Papers [Ame82, Smi81, Rom91]. Papyri [CWH^{+}52]. Paramagnetic [Pau31a]. parameter [Hat72, HCST73]. part [Sla70a, Seg77a, Seg77b]. Particle [TS52]. particles [Kos86, Sla85]. partisan [Tay69]. Pauling [Dus36, Par00]. PbS [MNS56]. Per-Olov [Hof67, CGLÖ76]. Percentages [WFS^{+}01]. Peredacha [Sla47]. Performing [Sla67d]. period [Sla73c]. Periodic [SS53, Sla37e, Sla49a, Sla53a, SK54, WFS^{+}01, Sla53f]. personal [Com56]. Perturbations [Sla27a]. Perturbed [Sla49a, Sla53f]. Phase [Kle70]. Phases [Ano37]. Phenomena [Sla25e]. Phenomenological [EAW15]. Philosophical [Sla56b, Smi81]. Philosophy [GHW09]. Photoconductivity [MNS56, Sla56a]. Photoelectron [Fis19]. Physical [MJW71, Sla29b, Sla49b]. Physically [Sla25d]. Physicist [Sla75c, Wil76]. Physicists [Hen03, Hen04, Hor00, Rid84]. Physics [BO71, Ehr75, EAW15, Gar70, GHW09, HBTW92, LP64, MMSW97, Pag34, SF33, Sla39, Sla46b, Sla48b, Sla49e, SF51, Sla55b, Sla58a, Sla65d, Sla67c, Sla69c, Sla70c, Sop80, Sop88, WP85, Wei70b, Sla49c, Sla68g, Sop76, Van64, Bar56, Coh56, Ett35a, Ett35b, Pag34, Tol56, Rom91]. Pisciculture [CWH^{+}52]. Plane [SS53, Sla53a, Sla64b, Sla67d, Woo62]. Plane-Wave [SS53]. Plants [CWH^{+}52]. Point [Sla40b]. Poisson [Don11, ZD11]. polarization [Sla50b]. Polarized [Sla68c]. Polyatomic [Sla31a, SJ72]. ponderable [Sla38a]. Position [SWS^{+}73, RFST85]. Possible [BU26]. Potential [SS53, Sla37e, Sla53a, SK54, Gás54, HMMT04, HM05]. Potentialfunktion [Gás54]. Potentials [Gás54, SWW69]. Power [Sla76b]. pp [OT80, Sla68b, T.75]. Practice [Par03]. Pragmatics [Hen04]. Premature [Kon98]. Present [Sla70b]. Press [Sla68b]. Pressure [Sla40b, Sla65d, Sla20]. Price [Sla68b, OT80, T.75]. Principles [LC10].

Prize [Ano67]. Probabilities [HS25, Sla25b, Sla71d]. Probability [Kos86]. Problem [SS53, Sla52a, Sla53a, Sla53g, SK54, Sla66, Sla68i, Sla70a]. Problems [Sla58a, Sla51f, Sla67a]. Proceedings [DP74, LW77, MJW71, SM73, Wel67, Löw67]. processing [DWZ22]. Professor [Jon76, Löw66a, PCR73, Sla71b]. Program [Sla56e, Sla56f]. Projected [Löw83]. propagator [AM06]. Properties [SKW62, RFST85]. Proposal [Sla56f]. prospects [Sla73a]. pseudospectral [Don11]. Publications [Ett35b].

QTP [Con10]. quality [RLER04]. Quanta [Sla28b]. Quantentheorie [BKS24b]. quantistica [Sla80]. Quantization [Sla68b]. Quantum [CGLÖ76, Cob72, DP74, Gar70, GHW09, Hen81, Hen03, Kon98, Löw66b, MMSW97, Par03, Par09, Pau31a, PW35, Sch90, Sla25d, Sla25e, Sla31d, Sla31e, Sla51j, Sla53j, Sla60c, Sla63c, Sla65a, Sla67c, Sla67g, Sla68g, Sla68h, Sla69c, SJ74, Sla74c, Sla75b, Sla77, Sla78, Sop76, Sop80, Sop88, Was81, vdW67, BKS24b, BKS24a, Boh67, Dus36, Sla73c, Ano53, Meg61, Mer69, Phi74, Sla64a, Sla80, Tay69, Bli69, CC61, Chi64, Cho74, CWH^{+}52, Dew63, Dus52, Hof67, Kle53, T.75, Wit52]. quest [Com56].

R [Sla68b]. R. [LP64]. Radiation [Sla24b, Sla25c, Sla27a, Sla27b, Sla51g, Was81, BKS24b, BKS24a, Boh67]. Radii [Sla64c]. radiovola [Sla47]. ray [Fis19]. reactivity [SM73]. Receives [Ano67]. Recordings [CDFS53a, CDFS53b, CDFS54]. Regarding [Sla65e]. Region [SS52]. relation [HMMT04]. Relations [Gou30]. relativistic [PAAC93]. Relativity [BU26]. Remarks [Sla51i, SWS^{+}73, SC76, Van77]. Replacement [UG25]. representation [BT88]. representations [RFST85]. Reproduction [PM12]. requirement [UG25]. Research [MJW71, Sla56b, Sla56e, Sla56f, Sla76a, SM73, Wel67]. Resonant [MS52a, SS52]. Results [Pau31a, Sla68i]. Review [Ano53, Bar56, Bli69, CC61, Chi64, Cho74, Coh56, CWH^{+}52, Dew63, Dus52, Ehr75, Ett35a, Fer48, Hay51, Hof67, Kle53, Kon80a, Kon80b, Meg61, Mer69, Pag34, Phi74, Sla56d, Sla56b, Sla64a, Sla67d, Sla68b, Sla68i, Sla75b, Sop78, Tay69, Tol56, Wit52]. Reviews [Gar70]. Revised [Sla56b]. Revisited [CS18]. Richtmeyer [Ett35a, Ett35b]. Road [Jan08]. Robert [Sla64d]. Role [Hen81]. rotating [GU26]. roteerende [GU26]. Rothery [Sla56b, Sla56d]. Row [WWS70]. Royal [Sla56b]. Rules [Fis19, Ree99, Tud93, VV91]. Rydberg [Sla28a].

S [CWH^{+}52, LP64]. sa [And77]. Sanibel [Löw67, Sla71b, Sla73b, Löw83, Löw85]. Scattering [RS36]. Schrödinger [Don11, Sla27b, ZD11]. Science [CGLÖ76, Hil15, Hoc90]. Sciences [Sla76a]. Scientific [Cob72, Ehr75, Löw85, Par03, Sla75d, Sop78]. scientifique [And77]. Screening [WFS^{+}01]. Second [MMS49, Sla68b]. Self [KS65, Phi74, Sla28d, Sla53c, Sla69a, SJ72, Sla72d, Sla74c, Sla78, Sla51i, Sla71c]. Self-Consistent [KS65, Sla53c, Sla69a, Sla72d, Sla74c, Sla78, Sla51i, Sla71c].

Self-Consistent-Field [SJ72]. Semiconducters [Sla67g]. Semiconductors [Mer69]. Semiempirical [Seg77a, Seg77b]. September [LW77, Sla49e]. series [Sla76b]. session [KMV77]. sets [RLER04]. Several [SWW69]. Shackle [CWH^{+}52]. Shaler [CWH^{+}52]. Shielding [Sla30a]. Shockley [Mol95]. short [Don11]. Simplification [Sla51e]. Simplified [KS54a, SK54, Don11]. Simulation [Fis19]. Simultaneous [Par00]. Slater [Ame82, Ano53, Bom01, Coh56, CWH^{+}52, Ett35a, Ett35b, Fer48, Kon80b, KMV77, Löw67, Meg61, Mer69, Mor82, OT80, Pag34, Phi74, Smi81, Sop78, T.75, Tay69, Wei70a, Amo97, AM06, And11b, And77, Ano67, Anoxx, Ano23, Bar56, Bli69, BG24, BT88, CS18, CC61, Chi64, Cho74, Con10, DWZ22, Dew63, Don11, Dus36, Dus52, Ehr75, ERW75, Fis19, Fol62, FP73, Gar70, Gou30, Har65, Hay51, Hen81, Hoc90, Hod08, Hor00, HMMT04, HM05, ISS06, Jan08, Jon76, Kle53, KU94, Kon80a, Kon83, Kon97, Kon98, Kos86, Kra09, LS70, Löw66a, Löw66b, Löw83, Mag14, Moh13, Nes10, Par00, PAAC93, Pra52, PM12, Ree99, RFST85, RH09, RLER04, Sch90, SHR13, Tal04, Tol56, Tud93, USOA13, Van76, Van77, VV91, Wag83, WFS^{+}01, Was81, Wei07, Wil76]. Slater [Wit52, ZD11, Hof67]. Slater-Type [Mag14, PM12, BT88]. Society [Sla56b, MJW71, Smi81]. Sodium [HS25]. Solid [HBTW92, Hof67, Löw66b, Löw67, Sla65e, Sla67e, Sla71a, Sla75d, Sla52d, Sla53h, Sla53i, Sop78, Ehr75]. Solid-State [Löw67, Sla67e, Sla71a, Sla75d, Sop78]. Solids [AFR68, Chi64, Cho74, Mer69, Phi74, RS36, Sla55a, Sla63c, Sla65a, Sla67g, SJ72, Sla74c, Sla78, Dew63, Sla51b, Sla53e, Sla53f, Sla54, Sla57, Sla67a, Sla68f, Sla68e, Sla72b, Sla75a, Sla76b, T.75, Phi74, Sla64a]. Soluble [Sla52a]. son [And77]. Sources [vdW67]. Space [Sla65b, RFST85, Sla70a]. spacings [Sla59c]. Spanish [SF51]. Spectra [Sla25a, Sla26d, Sla29a, Sla63a, UG26, GU26]. Spectroscopy [Fis19, Sla68b]. Speeches [CDFS53a, CDFS53b, CDFS54]. sphere [SC76]. spherically [Don11]. Spheroidal [SS52]. Spin [Löw83, Pau40, Sla68c]. Spin-Polarized [Sla68c]. Spin-Projected [Löw83]. Spinning [Sla26d, Tho26, UG26]. sponsorship [MJW71]. Squares [PM12]. SSMTG [Nes10]. State [And11b, Ehr75, HBTW92, Hof67, Löw66b, Löw67, Sla28c, Sla31d, Sla37b, Sla37c, Sla53h, Sla53i, Sla65e, Sla67e, Sla71a, Don11, PAAC93, Sla52d, Sla75d, ZD11, Sop78]. States [Sla68a, Cob72, Rid84]. Statistical [SW70, SW71, Sla72d, Sla72c, WWS70, Sla70b]. Statistics [Pau40]. status [Sla70b]. STO [PM12]. Strahlung [BKS24b]. Strength [MS52b, MS52a]. Structure [Ano37, CGLÖ76, CC61, Chi64, HS63, Joh65, Pau31a, Seg77a, Seg77b, Sla26d, Sla27c, Sla28d, Sla31c, Sla32b, Sla33b, Sla34b, Sla37g, Sla50b, Sla53d, Sla60c, Sla63b, Sla63c, Sla67d, UG26, Dew63, GU26, Sla53e, Sla53f, Sla54, Sla57, SM73, Meg61]. structuur [GU26]. Studies [WWS70]. Study [PAAC93, RH09, Wag83]. Suggestions [Sla65e]. Sulfide [Sla56a]. Sum [Moh13, AM06, HMMT04]. Summer [SM73]. Superconducting [Sla37b, Sla37c]. Superconductivity [BGH^{+}46]. Superconductors [MMS49]. Superlattices [Sla51d]. Surface [MMS49]. surfaces [Sla67b]. Survey [Löw85]. Susceptibility [Pau31a]. symbol [FP73]. symmetric [Don11, Sla70a]. Symmetry [SKW62, Sla65a, Sla65b, Sla72e]. Symposia [Löw85]. Symposium [Löw67, Sla71b, Sla73b, Wel67, Ano37]. System [SS52, Don11, ZD11]. Systematic [SHR13]. Systems [MMSW97, Sla25d].

Tables [KU94, Sla56b]. Tape [CDFS53a, CDFS53b, CDFS54]. Tc [Sla51a]. Teaching [WFS^{+}01]. Techniques [Seg77a]. Technology [Sla56f]. Temperature [Sla36b]. Temperatures [Sla49e]. Teoria [Sla80]. Teórica [SF51]. Terry [Sla67d]. TFD [Sla75a]. Thanks [Sla49d]. theorem [RLER04]. Theorems [Sla72a]. theoretic [RFST85]. Theoretical [EAW15, Ett35a, Ett35b, Hen03, Pag34, SF33, SF51, SM73]. Theorie [BG24]. Theory [Ano53, CC61, Chi64, Cho74, CWH^{+}52, Dus52, Ehr75, Hen81, Hof67, Kle53, Kon83, Kon98, Kra09, Löw66b, Löw67, MJW71, Mer69, Pau31a, Phi74, RS36, Sla25e, Sla27b, Sla29a, Sla31d, Sla31e, Sla33a, Sla37d, Sla41, Sla51j, Sla53b, Sla53j, Sla56a, Sla60c, Sla63c, Sla64a, Sla65a, Sla65e, Sla67e, Sla67g, Sla68d, Sla68h, Sla71a, Sla74c, Sla75d, Sla77, Sla78, Sla80, Sla87, Sop78, Tay69, Was81, Wit52, Bli69, BKS24b, BKS24a, Boh67, BG24, Con10, Dew63, Dus36, Kon97, Meg61, Sla51b, Sla53h, Sla53i, Sla56c, Sla59a, Sla67f, Sla68e, Sla72b, T.75, Sla75b]. Thermodynamical [Sla37h]. Thomas [MJW71, SK35]. Thoughtful [And11a]. Three [Pau31b]. Three-Electron [Pau31b]. tin [Sla74a]. Titanate [Sla50a]. Tool [WFS^{+}01]. Topics [BO71]. Total [SW70, SW71, Hat72, HCST73]. Traite [CWH^{+}52]. transistor [Hoc90]. Transition [Sla25b, Sla41, SMWW69, Sla71d, Sla87, WWS70]. Transmission [Cob72, Sla42, Sla59b, Sla60a]. Treatment [Sla71a]. Trends [MMSW97, WFS^{+}01]. Tribute [BO71, CGLÖ76, Hof67, LP64, Löw66b, PCR73]. Two [CS18, Jan08, SSK53]. Two-Electron [SSK53]. Type [Mag14, PM12, BT88]. Typescript [Sla75c].

U [BO71, Sla73b]. ulstrakorotkikh [Sla47]. Unconstrained [Moh13]. Undergraduate [Gar70]. Understanding [SHR13]. United [Cob72, Rid84]. units [DWZ22]. universal [Gás54]. Universelle [Gás54]. University [Ano37, Sla56b]. unmechanical [UG25]. unmechanischen [UG25]. Unrestricted [WP57]. USA [Ano37]. Using [Moh13, PM12, FP73].

Valence [Sla31a, Sla31b, Dus36]. Values [Hen03]. vanadium [Hat72, HCST73]. Vannevar [Gol92]. Vapor [HS25]. Verhaltens [UG25]. verification [BG24]. Versatile [WFS^{+}01]. versus [Mag14]. Very [Sla49e]. vi [OT80]. via [Woo62]. vie [And77]. view [Tay69]. viii [Sla68b]. Virial [Sla33b, Sla72a]. Virtual [Hen81, Kon98]. Vleck [Sla71b, Jan08]. Vol [Chi64, Cho74, Phi74, T.75]. Volume [CC61, Dew63, FK85, Mer69, Sla65a, Sla78]. vom [UG25]. vs [RFST85].

W [CWH^{+}52, Sla65d, SWS^{+}73]. Waals [CS18, SK31]. Wadham [LW77]. Warren [Sla56b, Wel67]. Wars [Sla67c, Sla68g]. Watson [MJW71]. Wave [BGH^{+}46, KS54b, KS54c, Pra52, PCR73, SS53, Sla28a, Sla28b, Sla32a, Sla37e, Sla53a, Sla65b, Sla67d, WP57, RH09, Sla29b, Sla46b, Sla51f, Sla70a, Sla64b, Woo62]. Wave-Function [Sla65b]. wave-functions [RH09]. Wave-Length [BGH^{+}46]. Waves [Sla37a, Sla58b, Sla85]. way [BG24]. Weg [BG24]. Welding [CWH^{+}52]. Wiley [OT80]. William [Sla56d, Sla56b, Mol95]. Wilson [CWH^{+}52]. Work [Sla53h, Sla53i, Was81]. Working [CWH^{+}52]. World [DP74]. Wulff [CWH^{+}52].

X [Con77, Fis19, Sla31c, Sla70b, Sla72a, SJ72, Sla73a, Sla74b, Sla75a]. X-ray [Fis19]. XI [Dus36]. xvii [T.75].

Years [Ehr75, Löw85, Par03, PCR73, Sop88]. York [MJW71, OT80, Sla56b, T.75]. Yorktown [MJW71]. Young [Sch90].

Zeeman [BU26]. Zones [Sla51d]. zur [BG24]. Zwang [UG25].

References

Alder:1968:EBS

[AFR68] Berni J. Alder, Sidney Fernbach, and Manuel Rotenberg, editors. Energy Bands of Solids, volume 8 of Methods in computational physics. Academic Press, New York, NY, USA, 1968. xvi + 300 pp. LCCN QA401 .M514 v.8.

Amovilli:2006:FPS

[AM06] Claudio Amovilli and Norman H. March. Feynman propagator and Slater sum for a model Hamiltonian motivated by H in an intense magnetic field. International Journal of Quantum Chemistry, 106 (2):533–541, ???? 2006. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

APS:1982:JCS

[Ame82] American Philosophical Society. John Clarke Slater Papers 1908– 1976. Document archive, 1982. URL http://www.amphilsoc.org/

           mole/view?docId=ead/Mss.B.SL2p-ead.xml. Contains roughly

50,000 items.

Amovilli:1997:EME

[Amo97] C. Amovilli. On the evaluation of the matrix elements between nonorthogonal Slater determinants. In McWeeny et al. [MMSW97], pages 343–348. ISBN 0-7923-4699-8. LCCN QD462 .Q348 1997.

Andre:1977:JCS

[And77] Jean-Marie André. John Clarke Slater, sa vie et son oeuvre scientifique. Revue des Questions Scientifiques, 148(4):445–456, October 1977. CODEN RQSCAN. ISSN 0035-2160. URL https://www.

rqs.be/app/views/revue.php?id=148.

Anderson:2011:MDN

[And11a] P. W. (Philip W.) Anderson. More and Different: Notes from a Thoughtful Curmudgeon. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4350-12-5 (hardcover), 981-4350-13-3 (paperback). ix + 412 pp. LCCN Q171 .A527 2011.

Anderson:2011:NMJ

[And11b] P. W. (Philip W.) Anderson. Nevill Mott, John Slater, and the “Magnetic state”. In More and Different: Notes from a Thoughtful Curmudgeon [And11a], pages 120–130. ISBN 981-4350-12-5 (hardcover), 981-4350-13-3 (paperback). LCCN Q171 .A527 2011.

Anonymous:1937:SSM

[Ano37] Anonymous, editor. Symposium on the Structure of Metallic Phases, Cornell University, Ithaca, NY, USA, 1–3 July 1937. Cornell University, Ithaca, NY, USA, 1937. LCCN ????

Anonymous:1953:BRJ

[Ano53] Anonymous. Book review: John C. Slater, Quantum Theory of Matter. Physics Today, 6(2):18–19, February 1953. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v6/
i2/p18_s3.

Anonymous:1967:JCS

[Ano67] Anonymous. John C. Slater receives 1967 Irving Langmuir Prize. Physics Today, 20(6):105, June 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v20/i6/p105_s1.

Anonymous:19xx:SJC

[Anoxx] Anonymous. Slater, John C. (John Clarke), 1900–1976. Social Networks and Archival Context Web site, 19xx. URL http://

n2t.net/ark:/99166/w6gm88vx;   https://snaccooperative.
org/view/38857821. SNAC ID 38857821.

Anonymous:2023:JCS

[Ano23] Anonymous. John C. Slater. AIP physics history network Web site.,

2023. URL https://history.aip.org/phn/11406001.html.

Barker:1956:BRB

[Bar56] E. F. Barker. Book review: Modern Physics, John C. Slater. American Journal of Physics, 24(4):289, April 1956. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v24/i4/p289_s1.

Bothe:1924:WEN

[BG24] W. Bothe and H. Geiger. Ein Weg zur experimentellen Nachprüfung der Theorie von Bohr, Kramers, und Slater. (German) [One way of experimental verification of the theory of Bohr, Kramers, and Slater]. Zeitschrift für Physik, 26(1):44, December 1924. CODEN ZEPYAA. ISSN 0044-3328.

Bitter:1946:SLC

[BGH^{+}46] F. Bitter, J. B. Garrison, J. Halpern, E. Maxwell, J. C. Slater, and C. F. Squire. Superconductivity of lead at 3-cm wave-length. Physical Review (2), 70(1–2):97–98, July 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v70/i1-2/p97_1.

Bohr:1924:QTR

[BKS24a] Niels Bohr, Henrik A. Kramers, and John Clarke Slater. The quantum theory of radiation. Philosophical Magazine, 47(281):785–802, May 1924. CODEN PHMAA4. ISSN 0031-8086. Reprinted in [Boh67], and in German translation in [BKS24b].

Bohr:1924:QSG

[BKS24b] Niels Bohr, Henrik A. Kramers, and John Clarke Slater. Über die Quantentheorie der Strahlung. (German) [On the quantum theory of radiation]. Zeitschrift für Physik, 24(1):69–87, December 1924. CODEN ZEPYAA. ISSN 0044-3328. URL http://link.

springer.com/article/10.1007/BF01327235.

Blinder:1969:BRB

[Bli69] S. M. Blinder. Book review: Quantum theory of matter, by John C. Slater. Journal of Chemical Education, 46(6):A447, July 1969. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

URL http://adsabs.harvard.edu/abs/1969JChEd..46..447B.

Brittin:1971:TMP

[BO71] Wesley Emil Brittin and Halis Odabasi, editors. Topics in Modern Physics: a Tribute to Edward U. Condon. Adam Hilger Ltd., Bristol, UK, 1971. ISBN 0-85274-195-2, 0-87081-010-3. xxxiii + 360 pp. LCCN QC71 .T63.

Bohr:1967:QTR

[Boh67] Niels Bohr. The quantum theory of radiation. In van der Waerden [vdW67], pages 159–176. LCCN QC174.12 S655.

Boman:2001:JCS

[Bom01] John Stewart Boman. John C. Slater (Clarke) (1900–76). In Cambridge Dictionary of American Biography, page ?? Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-40258-1. LCCN CT213 .C36 1995.

Braunss:1988:RST

[BT88] G. Braunss and E. Teubner. On the representation of Slater-type densities by Gaussian densities. Journal of Mathematical Physics, 29(8):1758–1760, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i8/p1758_s1.

Bichowsky:1926:PER

[BU26] F. R. Bichowsky and H. C. Urey. A possible explanation of the relativity doublets and anomalous Zeeman effect by means of a magnetic electron. Proceedings of the National Academy of Sciences of the United States of America, 12(2):80–85, February 15, 1926. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (elec-

tronic). URL http://www.jstor.org/stable/85075;   http:/
/www.pnas.org/content/12/2/80.full.pdf+html.

Chiu:1961:BRB

[CC61] Y. N. Chiu and L. Y. C. Chiu. Book review: Quantum Theory of Atomic Structure Volume I; Volume II, by John C. Slater. American Scientist, 49(3):332A–334A, September 1961. CODEN AM-SCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http:

//www.jstor.org/stable/27827894.

Condon:1953:TRIa

[CDFS53a] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 21(2):150, February 1953. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/21/150/2;   http://link.aip.org/
link/ajpias/v21/i2/p150/s2.

Condon:1953:TRIb

[CDFS53b] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 21(3):227, March 1953. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/21/227/1;   http://link.aip.org/
link/ajpias/v21/i3/p227/s1.

Condon:1954:TRIc

[CDFS54] Edward U. Condon, K. K. Darrow, Enrico Fermi, and John C. Slater. The tape recordings of important speeches. American Journal of Physics, 22(4):246, April 1954. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/?AJP/22/246/2;   http://link.aip.org/link/
ajpias/v22/i4/p246/s2.

Calais:1976:QSM

[CGLÖ76] Jean-Louis Calais, Osvaldo Goscinski, Jan Linderberg, and Yngve Öhrn. Quantum Science: Methods and Structure: a Tribute to Per-Olov Löwdin. Plenum Press, New York, NY, USA; London, UK, 1976. ISBN 0-306-30968-8. ix + 595 pp. LCCN QC174.125 .Q371.

Chiu:1964:BRB

[Chi64] L. Y. C. Chiu. Book review: Quantum Theory of Molecules and Solids, Vol. 1: Electronic Structure of Molecules, by John C. Slater. American Scientist, 52(1):86A–87A, March 1964. CODEN AM-SCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http:

//www.jstor.org/stable/27838941.

Chong:1974:BRB

[Cho74] D. P. Chong. Book review: Quantum Theory of Molecules and Solids, Vol. 4, by John C. Slater. American Scientist, 62(6):725, November 1974. CODEN AMSCAC. ISSN 0003-0996 (print),

1545-2786 (electronic). URL http://www.jstor.org/stable/
27845185.

Coben:1972:SET

[Cob72] Stanley Coben. The scientific establishment and the transmission of quantum mechanics to the United States, 1919–32. The American Historical Review, 76(2):442–466, April 1972. ISSN 0002-8762 (print), 1937-5239 (electronic). URL http://www.jstor.

org/stable/1858707.

Cohen:1956:BRJ

[Coh56] E. Richard Cohen. Book review: John C. Slater, Modern Physics. Physics Today, 9(6):22, June 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v9/i6/p22_s1.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Connolly:1977:XM

[Con77] John W. D. Connolly. The Xα method. In Segal [Seg77a], chapter 4, pages 105–132. ISBN 0-306-33507-7. ISSN 0146-0420. LCCN QD461 .S396.

Connolly:2010:QJC

[Con10] John W. D. Connolly. QTP in the 60s: John C. Slater and the beginnings of density functional theory. Molecular Physics, 108(21– 23):2863–2866, 2010. CODEN MOPHAM. ISSN 0026-8976 (print),

1362-3028 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00268976.2010.496744. Special Issue: Electrons,

Molecules, Solids, and Biosystems: Fifty Years of the Quantum Theory Project. See [Sla51e, Gás54, KS65, Sla72d, Con77].

Cances:2018:VWI

[CS18] Eric Cancès and L. Ridgway Scott. Van der Waals interactions between two hydrogen atoms: The Slater–Kirkwood method revisited. SIAM Journal on Mathematical Analysis, 50(1):381–410, ???? 2018. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crawford:1952:BRBb

[CWH^{+}52] J. H. Crawford, George N. Wolcott, D. Hey, Michael B. Shimkin, Ralph C. Corley, Joshua H. Rosenbloom, E. C. Miller, Marjorie B.

Snyder, Frank T. Knapp, and Norman Taylor. Book review: Quantum Theory of Matter by John C. Slater; The Injurious Insects of the British Commonwealth by J. W. Evans; Traite de Pisciculture by Marcel Huet; The Old Egyptian Medical Papyri by Chauncey D. Leake; Claude Bernard and the Experimental Method in Medicine by J. M. D. Olmsted and E. Harris Olmsted; Dynamic Aspects of Biochemistry by Earnest Baldwin; Mathematics at the Fireside by G. L. S. Shackle; Metallurgy for Engineers: Casting, Welding, and Working by John Wulff, Howard F. Taylor and Amos J. Shaler; My Brother, My Enemy by Mitchell Wilson; Freshwater Fishery Biology by Karl F. Lagler; Plants of the Bible by Harold N. Moldenke and Alma L. Moldenke . The Scientific Monthly, 75(5):314–319, November 1952. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-

7513 (electronic). URL http://www.jstor.org/stable/20784.

Dewar:1963:BRB

[Dew63] Michael J. S. Dewar. Book review: Quantum theory of molecules and solids. Volume 1, electronic structure of molecules, by John C. Slater. Journal of Chemical Education, 40(11):615, November 1963. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

URL http://adsabs.harvard.edu/abs/1963JChEd..40..615D.

Demos:1952:MLE

[DKS52] P. T. Demos, A. F. Kip, and John C. Slater. The M.I.T. Linear Electron Accelerator. Journal of Applied Physics, 23(1):53–65, January 1, 1952. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://link.aip.org/link/

japiau/v23/i1/p53/s1.

Dong:2011:SNS

[Don11] Xuanchun Dong. A short note on simplified pseudospectral methods for computing ground state and dynamics of spherically symmetric Schrödinger–Poisson–Slater system. Journal of Computational Physics, 230(22):7917–7922, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004542.

Daudel:1974:WQC

[DP74] Raymond Daudel and Bernard Pullman, editors. The World of Quantum Chemistry. Proceedings of the First International Congress of Quantum Chemistry Held at Menton, France, July 4– 10, 1973. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0421-X. LCCN QD462.A1 I57 1973.

Dushman:1936:EQT

[Dus36] Saul Dushman. Elements of the quantum theory. XI. Slater– Pauling theory of valence bonds. Journal of Chemical Education, 13(8):385–393, August 1936. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Dushman:1952:BRB

[Dus52] Saul Dushman. Book review: Quantum Theory of Matter by John C. Slater. Journal of Chemical Education, 29(4):207, April 1952. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

URL http://adsabs.harvard.edu/abs/1952JChEd..29..207D.

Dang:2022:NES

[DWZ22] Duy-Khoi Dang, Leighton W. Wilson, and Paul M. Zimmerman. The numerical evaluation of Slater integrals on graphics processing units. Journal of Computational Chemistry, 43(25):1680–1689, September 30, 2022. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The Physics of Ettore Majorana: Phenomenological, Theoretical, and Mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Ehrenreicii:1975:BRF

[Ehr75] Henry Ehrenreicii. Book review: Fifty years in physics and chemistry: Solid State and Molecular Theory. A Scientific Biography, by John C. Slater. Science, 188(4190):839–840, May 23, 1975. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/1740606.

Ellis:1975:ADS

[ERW75] D. E. Ellis, A. Rosén, and P. F. Walch. Applications of the Dirac– Slater model to molecules. International Journal of Quantum Chemistry, 9(S9):351–358, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Ettlinger:1935:BRB

[Ett35a] H. J. Ettlinger. Book review: Introduction to Theoretical Physics by F. K. Richtmeyer, J. C. Slater, and N. H. Frank. American Mathematical Monthly, 42(6):389–390, June/July 1935. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2301365.

Ettlinger:1935:RPI

[Ett35b] H. J. Ettlinger. Recent publications: Introduction to Theoretical Physics, by F. K. Richtmeyer, J. C. Slater, and N. H. Frank. American Mathematical Monthly, 42(6):389–390, June/July 1935. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ferraro:1948:BRB

[Fer48] V. C. A. Ferraro. Book review: Mechanics by J. C. Slater and N. H. Frank. Mathematical Gazette, 32(302):317, December 1948. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/stable/3609914.

Fisher:2019:ICM

[Fis19] Aidan A. E. Fisher. An introduction to coding with Matlab: Simulation of X-ray photoelectron spectroscopy by employing Slater’s rules. Journal of Chemical Education, 96(7):1502–1505, June 2019. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a Centenary Volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Foldy:1962:AFS

[Fol62] Leslie L. Foldy. Antisymmetric functions and Slater determinants. Journal of Mathematical Physics, 3(3):531–539, March 1962. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/
i3/p531_s1.

FroeseFischer:1973:EES

[FP73] C. Froese Fischer and D. W. B. Prentice. Exact evaluation of Slater integrals using ALTRAN, a symbol manipulation language. Computer Physics Communications, 6(4):157–164, October 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900970.

Garber:1970:BFR

[Gar70] Elizabeth Garber. Book and film reviews: For the undergraduate physics major: Concepts and Development of Quantum Mechanics, by John C. Slater. The Physics Teacher, 8(9):535–536, December 1970. CODEN PHTEAH. ISSN 0031-921X (print),

1943-4928 (electronic). URL http://tpt.aapt.org/resource/1/
phteah/v8/i9/p535_s2.

Gaspar:1954:AHF

[Gás54] R. Gáspár. Über eine Approximation des Hartree–Fockschen Potentials Durch eine Universelle Potentialfunktion. (German) [On an approximation of the Hartree–Fock potential by a universal potential function]. Acta physica Academiae Scientiarum Hungaricae, 3(3–4):263–286, April 1954. CODEN APAHAQ.

ISSN 0001-6705. URL http://www.springerlink.com/content/

w24337x7040846x5/. In this paper, the Hartree–Fock–Slater X-α equations [Sla51e, Sla72d] are rederived by an independent method, with the prediction of a value α = 2/3, instead of Slater’s original α = 1. The same factor was later rediscovered in independent work [KS65].

Greenberger:2009:CQP

[GHW09] Daniel M. Greenberger, Klaus Hentschel, and Friedel Weinert, editors. Compendium of Quantum Physics: Concepts, Experiments, History and Philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-70626-7, 3-540-70622-4. xvi + 901 pp. LCCN C174.12 .C66 2009.

Goldberg:1992:ICO

[Gol92] Stanley Goldberg. Inventing a climate of opinion: Vannevar Bush and the decision to build the bomb. Isis, 83(3):429–452, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233904.

Goudsmit:1930:EHS

[Gou30] S. Goudsmit. An extension of Houston’s and Slater’s multiplet relations. Physical Review, 35(11):1325–1336, June 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Goudsmit:1926:RES

[GU26] S. Goudsmit and G. E. Uhlenbeck. Over het roteerende electron en de structuur der spectra. (Dutch) [On the rotating electron and the structure of the spectra]. Physica, 6(??):273–290, ???? 1926. CO-DEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Harris:1965:DFS

[Har65] Frank E. Harris. Discussion following Slater’s paper. Journal of Chemical Physics, 43(10):S17, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). See [Sla65c].

Hattox:1972:CMT

[Hat72] Thomas Maxwell Hattox. A calculation of the magnetization and total energy of vanadium as a function of lattice parameter. Ph.D. thesis, Department of Physics, University of Florida, Gainesville, FL, USA, 1972. ix + 92 pp.

Haynes:1951:BRB

[Hay51] Sherwood K. Haynes. Book review: Microwave Electronics by John C. Slater. American Journal of Physics, 19(2):133, February 1951. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v19/i2/p133_
s2.

Hoddeson:1992:CMC

[HBTW92] Lillian Hoddeson, Ernest Brown, Jürgen Teichmann, and Spencer Weart, editors. Out of the Crystal Maze: Chapters from the History of Solid State Physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-505329-X. xx + 697 pp. LCCN QC176 .O98 1992. URL http://catdir.loc.

gov/catdir/enhancements/fy0603/89033498-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy0603/89033498-t.
html;   http://www.gbv.de/dms/bowker/toc/9780195053296.
pdf.

Hattox:1973:CMT

[HCST73] Thomas M. Hattox, James B. Conklin, Jr., John C. Slater, and Samuel B. Trickey. Calculation of the magnetization and total energy of vanadium as a function of lattice parameter. The Journal of Physics and Chemistry of Solids, 34(10):1627–1638, October 1973. CODEN JPCSAW. ISSN 0022-3697 (print), 1879-2553 (electronic).

URL http://adsabs.harvard.edu/abs/1973JPCS...34.1627H.

Hendry:1981:BKS

[Hen81] John Hendry. Bohr–Kramers–Slater: a virtual theory of virtual oscillators and its role in the history of quantum mechanics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 25(2):189–221, July 1981. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Hendry:2003:AET

[Hen03] Robin Findlay Hendry. Autonomy, explanation, and theoretical values: Physicists and chemists on molecular quantum mechanics. Annals of the New York Academy of Sciences, 988(1):44–58, May 2003. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Hendry:2004:PCP

[Hen04] Robin Findlay Hendry. The physicists, the chemists, and the pragmatics of explanation. Philosophy of Science, 71(5):1048–1059, 2004. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/3693328.

Hiltzik:2015:BSE

[Hil15] Michael A. Hiltzik. Big Science: Ernest Lawrence, the Cyclotron, and the Birth of the Military–Industrial Complex. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-7575-5 (hardcover), 1-4516-7576-3 (paperback), 1-4516-7603-4 (e-book). x + 512 pp. LCCN QC787.C8 H55 2015.

Howard:2005:SNE

[HM05] I. A. Howard and N. H. March. Slater’s nonlocal exchange potential and beyond. International Journal of Quantum Chemistry, 102(1): 64–71, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Howard:2004:SSO

[HMMT04] I. A. Howard, A. Minguzzi, N. H. March, and M. P. Tosi. Slater sum for the one-dimensional sech^{2} potential in relation to the kinetic energy density. Journal of Mathematical Physics, 45(6):2411–2419, June 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i6/p2411_s1.

Hoch:1990:JCS

[Hoc90] P. K. Hoch. John Clarke Slater and the science for the transistor. European Journal of Physics, 11(5):283–291, September 1, 1990. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL http://iopscience.iop.org/0143-0807/11/5/005/.

Hoddeson:2008:SJC

[Hod08] Lillian Hoddeson. Slater, John Clarke. In Charles Coulston Gillispie, Frederic Lawrence Holmes, and Noretta Koertge, editors, Complete Dictionary of Scientific Biography, page ?? Gale Virtual Reference Library, Farmington Hills, MI, USA, 2008. ISBN 0-684-31559-9 (e-book). LCCN ???? URL http://www.encyclopedia.

com/doc/1G2-2830905335.html.

Hoffman:1967:BRB

[Hof67] Joseph G. Hoffman. Book review: Quantum Theory of Atoms, Molecules and the Solid State: A Tribute to John C. Slater, by Per-Olov Löwdin. Physics Today, 20(11):91–92, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/20/91/1;
 http://www.physicstoday.org/resource/1/phtoad/v20/i11/
p91_s1.

Hornbuckle:2000:SJC

[Hor00] Adam R. Hornbuckle. Slater, John Clarke (1900–1976), physicists. In American National Biography Online, 2000–02, page ?? Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-860669-9. LCCN ???? URL http://www.anb.org/

articles/13/13-02041.html.

Harrison:1925:LBA

[HS25] George R. Harrison and J. C. Slater. Line breadths and absorption probabilities in sodium vapor. Physical Review (2), 26(2):176– 188, August 1, 1925. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v26/i2/p176_1.

Herman:1963:ASC

[HS63] Frank Herman and Sherwood Skillman. Atomic Structure Calculations. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963. various pp. LCCN QC174.45 .H43.

Ixaru:2006:FCS

[ISS06] L. Gr. Ixaru, N.-S. Scott, and M. P. Scott. Fast computation of the Slater integrals. SIAM Journal on Scientific Computing, 28 (4):1252–1274, January 2006. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Janssen:2008:VVS

[Jan08] Michel Janssen. Van Vleck and Slater: Two Americans on the road to matrix mechanics. In American Physical Society, 2008 APS April Meeting and HEDP/HEDLA Meeting, April 11–15, 2008, page ?? American Physical Society, Ridge, NY 11961, USA, April 2008. URL http:/

/adsabs.harvard.edu/abs/2008APS..APR.R6001J;   http://
www.stolaf.edu/depts/physics/colloquium/Janssen_10-07-

09.html. Abstract #R6.001.

Johnson:1965:EBS

[Joh65] Keith Huber Johnson. The Electronic Band Structure of Ordered Beta-Brass by the Green’s Function Method. Ph.D. Thesis, Temple University, Philadelphia, PA, USA, 1965. 174 pp. URL http://

search.proquest.com/docview/302144664.

Jones:1976:OPJ

[Jon76] H. Jones. Obituary: Professor J. C. Slater. Nature, 263(5580):804, October 28, 1976. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v263/n5580/pdf/263804a0.pdf.

Kaiser:2010:BMD

[Kai10] David Kaiser, editor. Becoming MIT: moments of decision. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-11323-6 (hardcover), 0-262-51815-5 (paperback), 0-262-28953-9 (e-book). vi + 207 pp. LCCN T171.M428 B43 2010. URL http://www.jstor.

org/stable/10.2307/j.ctt5hhbpp.

Klein:1953:BRBb

[Kle53] Martin J. Klein. Book review: Quantum Theory of Matter by John C. Slater. Philosophy of Science, 20(4):344–345, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185048.

Klein:1970:FPB

[Kle70] Martin J. Klein. The first phase of the Bohr–Einstein dialogue. Historical Studies in the Physical Sciences, 2(??):iv + 1–39, ???? 1970. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757302.

Koster:1977:MSJ

[KMV77] George F. Koster, Philip McCord Morse, and J. H. (John Hasbrouck) Van Vleck. Memorial session for John C. Slater / 1977 February 7. Sound recording, 1977. URL http:

//libserv.aip.org:81/ipac20/ipac.jsp?uri=full=3100001~!
25813!0&profile=newcustom-icos.

Koningstein:1980:BRBb

[Kon80a] J. A. Koningstein. Book review: The Calculation of Molecular Orbitals, by John C. Slater. American Scientist, 68(3):321, May 1980. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27849853.

Koningstein:1980:BRJ

[Kon80b] J. A. Koningstein. Book review: John C. Slater, The Calculation of Molecular Orbitals. American Scientist, 68(3):321, May 1980. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27849853.

Konno:1983:SEG

[Kon83] Hiroyuki Konno. Slater’s evidence for the genesis of the Bohr– Kramers–Slater theory. Historia Scientiarum = International journal of the History of Science Society of Japan, 25(??):39–52, September 1983. CODEN HISCDU. ISSN 0285-4821. URL https:

//ci.nii.ac.jp/naid/110009837441.

Konno:1997:JCS

[Kon97] Hiroyuki Konno. J. C. Slater’s optical theory in 1925: another dispersion theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(1):59–64, June 1997. CODEN HISCDU. ISSN 0285-4821.

Konno:1998:STL

[Kon98] Hiroyuki Konno. Slater’s theory of line breadth and the virtual oscillator model: A premature attempt in the old quantum theory. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 7(3):213–226, March 1998. CODEN HISCDU. ISSN 0285-4821.

Kostin:1986:PDB

[Kos86] M. D. Kostin. Probability density for Bose–Einstein and Fermi– Dirac particles: Slater–Kahn functions. Journal of Mathematical Physics, 27(12):3014–3015, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kragh:2009:BKS

[Kra09] Helge Kragh. Bohr–Kramers–Slater theory. In Greenberger et al. [GHW09], pages 62–64. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Koster:1954:SIC

[KS54a] G. F. Koster and J. C. Slater. Simplified impurity calculation. Physical Review (2), 96(5):1208–1223, December 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v96/i5/p1208_1.

Koster:1954:WFIa

[KS54b] G. F. Koster and J. C. Slater. Wave functions for impurity levels. Physical Review (2), 94(5):1392, June 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v94/i5/p1392_1.

Koster:1954:WFIb

[KS54c] G. F. Koster and J. C. Slater. Wave functions for impurity levels. Physical Review (2), 95(5):1167–1176, September 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v95/i5/p1167_1.

Kohn:1965:SCE

[KS65] W. Kohn and L. J. Sham. Self-consistent equations including exchange and correlation effects. Physical Review (2), 140(4A): A1133–A1138, November 1965. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.140.A1133;   http://www.nobelprize.
org/nobel_prizes/chemistry/laureates/1998/. The authors of

this paper rediscover an earlier result [Gás54], and obtain the same value α = 2/3, compared to Slater’s original α = 1 [Sla51e, Sla72d].

Kollar:1994:GFS

[KU94] J. Kollár and B. Ujfalussy. General formulas for the Slater– Koster tables. Journal of Computational Physics, 110(1):187– 189, January 1994. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199918471014X.

Lalena:2010:PIM

[LC10] John N. Lalena and David A. Cleary. Principles of Inorganic Materials Design. Wiley, New York, NY, USA, second edition, 2010. ISBN 0-470-40403-5 (hardcover). xxvi + 585 pp. LCCN QD151.3

.L35 2010. URL http://onlinelibrary.wiley.com/book/10.
1002/9780470567548.

Lowdin:1966:CPJ

[Löw66a] Per-Olov Löwdin. Comments on Professor John C. Slater’s paper, “Cohesion in Monovalent Metals”. In Quantum Theory of Atoms, Molecules, and the Solid State: a Tribute to John C. Slater [Löw66b], pages 15–?? LCCN QC174.1 .L6. See [Sla30b].

Lowdin:1966:QTA

[Löw66b] Per-Olov Löwdin, editor. Quantum Theory of Atoms, Molecules, and the Solid State: a Tribute to John C. Slater. Academic Press, New York, NY, USA, 1966. xvi + 641 pp. LCCN QC174.1 .L6.

Lowdin:1967:PIS

[Löw67] Per-Olov Löwdin, editor. Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, held in honor of John Clarke Slater at Sanibel Island, Florida, January 16–20, 1967, volume 1 of International Journal of Quantum Chemistry. Symposium. Interscience Publishers, New York, NY, USA, 1967. LCCN QD 462 A1 I58 1967.

Lowdin:1983:SCE

[Löw83] Per-Olov Löwdin. On the Sanibel coefficients in the expansion of spin-projected Slater determinants. International Journal of Quantum Chemistry, 24(6):729–745, ???? 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1985:YSS

[Löw85] P.-O. Löwdin. 25 years of Sanibel Symposia — a brief historic and scientific survey. International Journal of Quantum Chemistry.

Quantum Chemistry Symposium, 19:19–37, 1985. CODEN IJQSDI. ISSN 0161-3642.

Lowdin:1964:MOC

[LP64] Per-Olov Löwdin and Bernard Pullman, editors. Molecular Orbitals in Chemistry, Physics, and Biology: a Tribute to R. S. Mulliken. Academic Press, New York, NY, USA, 1964. xiii + 578 pp. LCCN QD461 .L872 1964.

Larson:1970:DMA

[LS70] Everett G. Larson and Darwin W. Smith. A density matrix analysis of Slater’s hyper-Hartree–Fock method. International Journal of Quantum Chemistry, 5(S4):87–94, January 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowde:1977:IEM

[LW77] R. D. Lowde and E. P. Wohlfarth, editors. Itinerant-electron Magnetism: Proceedings of the International Conference Held at Wadham College, Oxford, September 12–15, 1976. North-Holland Pub. Co., Amsterdam, The Netherlands, 1977. ISBN 0-444-85000-7 (hardcover). LCCN QC754.2.B35 I84.

Magalhaes:2014:GTO

[Mag14] Alexandre L. Magalhães. Gaussian-type orbitals versus Slater-type orbitals: a comparison. Journal of Chemical Education, 91(12): 2124–2127, September 2014. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Martin:2015:WNC

[Mar15] Joseph D. Martin. What’s in a name change? Physics in Perspective (PIP), 17(1):3–32, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-014-0151-7.

Meggers:1961:BRJ

[Meg61] William F. Meggers. Book review: John C. Slater, Quantum theory of atomic structure. Physics Today, 14(5):48–50, May 1961. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v14/
i5/p48_s1.

Mermin:1969:BRJ

[Mer69] N. D. Mermin. Book review: John C. Slater, Quantum Theory of Molecules and Solids, Volume 3: Insulators, Semiconductors, and Metals. Physics Today, 22(7):89–91, July 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v22/i7/p89_s1.

Marcus:1971:CMB

[MJW71] P. M. Marcus, J. F. Janak, and A. R. Williams, editors. Computational Methods in Band Theory: Proceedings of a Conference held at the IBM Thomas J. Watson Research Center, Yorktown Heights, New York, May 14–15, 1970, under the joint sponsorship of IBM and the American Physical Society, The IBM Research Symposia Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 1-4684-1892-0 (print),

1-4684-1890-4 (e-book). URL http://link.springer.com/book/
10.1007/978-1-4684-1890-3.

Maxwell:1949:SIN

[MMS49] E. Maxwell, P. M. Marcus, and J. C. Slater. Surface impedance of normal and superconductors at 24,000 megacycles per second. Physical Review (2), 76(9):1332–1347, November 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v76/i9/p1332_1.

McWeeny:1997:QSC

[MMSW97] Roy McWeeny, Jean Maruani, Yves G. Smeyers, and Stephen Wilson, editors. Quantum Systems in Chemistry and Physics: Trends in Methods and Applications, volume 16 of Topics in molecular organization and engineering. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4699-8. LCCN QD462 .Q348 1997.

Mahlman:1956:PP

[MNS56] G. W. Mahlman, W. B. Nottingham, and John C. Slater. Photoconductivity of PbS. In R. G. Breckenridge, B. R. Russell, and E. E. Hahn, editors, Photoconductivity Conference: Held at Atlantic City. November 4–6, 1954, pages 489–508. Wiley, New York, NY, USA, 1956. LCCN QC612.P5 P5 1954.

Mohlenkamp:2013:CIC

[Moh13] Martin J. Mohlenkamp. Capturing the interelectron cusp using a geminal layer on an unconstrained sum of Slater determinants. SIAM Journal on Applied Mathematics, 72(6):1742–1771, ???? 2013. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Moll:1995:WBS

[Mol95] John L. Moll. William Bradford Shockley: February 13, 1910– August 12, 1989. Biographical memoirs — National Academy of Sciences of the United States of America, 68:304–325, 1995. CO-DEN BMNSAC. ISBN 0-309-56440-9. ISSN 0077-2933. URL

http://www.nap.edu/catalog.php?record_id=4990;   http:/
/www.nap.edu/catalog/4990.html.

Morse:1982:JCS

[Mor82] Philip McCord Morse. John Clarke Slater: December 22, 1900– July 25, 1976: a biographical memoir. National Academy of Sciences, 53:296–321, 1982. URL http://www.nap.edu/html/

biomems/jslater.pdf.

Maier:1952:FSM

[MS52a] L. C. Maier and J. C. Slater. Field strength measurements in resonant cavities. Journal of Applied Physics, 23(1):68–77, January 1, 1952. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://link.aip.org/link/

japiau/v23/i1/p68/s1.

Maier:1952:DFS

[MS52b] L. C. Maier and John C. Slater. Determination of field strength in a linear accelerator cavity. Journal of Applied Physics, 23(1):78– 83, January 1, 1952. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://link.aip.org/

link/japiau/v23/i1/p78/s1.

Nesbet:2010:JCS

[Nes10] Robert K. Nesbet. John C. Slater and the SSMTG in 1954– 1956. Molecular Physics, 108(21–23):2867–2869, 2010. CO-DEN MOPHAM. ISSN 0026-8976 (print), 1362-3028 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/

00268976.2010.515256. Special Issue: Electrons, Molecules, Solids, and Biosystems: Fifty Years of the Quantum Theory Project.

Orville-Thomas:1980:BCM

[OT80] W. Orville-Thomas. The calculation of molecular orbitals, John C. Slater, John Wiley, New York, 1979, pp. vi + 104, price £11.50. Journal of Molecular Structure, 64:300, May 1980. CO-DEN JMOSB4. ISSN 0022-2860 (print), 1872-8014 (electronic).

URL http://adsabs.harvard.edu/abs/1980JMoSt..64R.300O;
 http://www.sciencedirect.com/science/article/pii/0022286080801529.

Pisani:1993:SRE

[PAAC93] Lorenzo Pisani, Jean-Marie Andre, Marie-Claude Andre, and Enrico Clementi. Study of relativistic effects in atoms and molecules by the kinetically balanced LCAO approach: Ground state of hydrogen and of hydrogenic atoms in Slater and Gaussian basis functions. Journal of Chemical Education, 70(11):894, November 1993. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Page:1934:BRT

[Pag34] Leigh Page. Book review: Theoretical physics: Introduction to Theoretical Physics by J. C. Slater and N. H. Frank. Science, 80 (2086):590, December 21, 1934. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1660022.

Park:2000:CSD

[Par00] B. S. Park. The contexts of simultaneous discovery: Slater, Pauling, and the origins of hybridisation. Studies in History and Philosophy of Modern Physics, 31(4):451–474, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000277.

Park:2003:HQC

[Par03] Buhm Soon Park. The ‘hyperbola of quantum chemistry’: the changing practice and identity of a scientific discipline in the early years of electronic digital computers, 1945–65. Annals of Science, 60(3):219–247, 2003. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.ingentaconnect.com/
content/tandf/tasc/2003/00000060/00000003/art00001.

Park:2009:BAM

[Par09] Buhm Soon Park. Between accuracy and manageability: Computational imperatives in quantum chemistry. Historical Studies in the Natural Sciences, 39(1):32–62, February 2009. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Pauling:1931:NCBa

[Pau31a] Linus Pauling. The nature of the chemical bond. application of results obtained from the quantum mechanics and from a theory of paramagnetic susceptibility to the structure of molecules. Journal of the American Chemical Society, 53(4):1367–1400, April 6, 1931. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984. See also [Pau31b, Sla31a].

Pauling:1931:NCBb

[Pau31b] Linus Pauling. The nature of the chemical bond. II. The oneelectron bond and the three-electron bond. Journal of the American Chemical Society, 53(9):3225–3237, September 5, 1931. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984. See also [Pau31a, Sla31a].

Pauli:1940:CBS

[Pau40] W. Pauli. The connection between spin and statistics. Physical Review (2), 58(8):716–722, October 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v58/i8/p716_1.

Price:1973:WMF

[PCR73] William Charles Price, Seymour S. Chissick, and Tom Ravensdale, editors. Wave Mechanics: the First Fifty Years. A Tribute to Professor Louis De Broglie, Nobel Laureate, on the 50th Anniversary of the Discovery of the Wave Nature of the Electron. Wiley, New York, NY, USA, 1973. ISBN 0-470-69731-8. 435 pp. LCCN QC174.22 .W38 1973.

Phillips:1974:BRJ

[Phi74] James C. Phillips. Book review: John C. Slater, Quantum Theory of Molecules and Solids Vol. 4: The Self Consistent Field for Molecules and Solids. Physics Today, 27(12):49–50, December 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v27/i12/p49_s1.

Pye:2012:LSF

[PM12] Cory C. Pye and Colin J. Mercer. On the least-squares fitting of Slater-type orbitals with Gaussians: Reproduction of the STO– NG fits using Microsoft Excel and Maple. Journal of Chemical Education, 89(11):1405–1410, September 2012. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Pratt:1952:WFE

[Pra52] George W. Pratt, Jr. Wave functions and energy levels for Cu^{+} as found by the Slater approximation to the Hartree–Fock equations. Physical Review (2), 88(6):1217–1224, December 15, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v88/i6/p1217_1.

Pauling:1935:IQM

[PW35] Linus Pauling and E. Bright Wilson, Jr. Introduction to Quantum Mechanics: With Applications to Chemistry. McGraw-Hill, New York, NY, USA, 1935. xi + 468 pp. LCCN QC174.1 .P34. Reprinted by Dover, 1985.

Reed:1999:GSR

[Ree99] James L. Reed. The genius of Slater’s rules. Journal of Chemical Education, 76(6):802, June 1999. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Regier:1985:GVS

[RFST85] Philip E. Regier, Jacob Fisher, B. S. Sharma, and Ajit J. Thakkar. Gaussian vs. Slater representations of d orbitals: An information theoretic appraisal based on both position and momentum space properties. International Journal of Quantum Chemistry, 28(4): 429–449, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reinhardt:2009:CDW

[RH09] Peter Reinhardt and Philip E. Hoggan. Cusps and derivatives for wave-functions expanded in Slater orbitals: A density study. International Journal of Quantum Chemistry, 109(14):3191–3198, ???? 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rider:1984:AOE

[Rid84] Robin E. Rider. Alarm and opportunity: Emigration of mathematicians and physicists to Britain and the United States, 1933– 1945. Historical Studies in the Physical Sciences, 15(1):107–176, ???? 1984. CODEN HSPSAS. ISSN 0073-2672. URL http://

www.jstor.org/stable/27757544.

Rico:2004:AET

[RLER04] J. Fernández Rico, R. López, I. Ema, and G. Ramírez. Accuracy of the electrostatic theorem for high-quality Slater and Gaussian basis sets. International Journal of Quantum Chemistry, 100(2): 221–230, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Romer:1991:EMP

[Rom91] Robert H. Romer. Editorial: Memorable papers from the American Journal of Physics, 1933–1990. American Journal of Physics, 59(3):201–207, March 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Rudberg:1936:TIS

[RS36] E. Rudberg and J. C. Slater. Theory of inelastic scattering of electrons from solids. Physical Review (2), 50(2):150–158, July 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v50/i2/
p150_1.

Slater:1976:ROS

[SC76] John C. Slater and John W. D. Connolly. Remarks on the overlapping-sphere method for molecular orbitals. International Journal of Quantum Chemistry, 10(S10):141–146, January 19–24, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Schweber:1990:YJC

[Sch90] S. S. Schweber. The young John Clarke Slater and the development of quantum chemistry. Historical Studies in the Physical and Biological Sciences, 20(2):339–406, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://www.

jstor.org/stable/27757647.

Segal:1977:SMEa

[Seg77a] Gerald A. Segal, editor. Semiempirical Methods of Electronic Structure Calculation. Part A. Techniques, volume 7 of Modern theoretical chemistry. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-33507-7. ISSN 0146-0420. xvii + 274 pp. LCCN QD461 .S396.

Segal:1977:SMEb

[Seg77b] Gerald A. Segal, editor. Semiempirical Methods of Electronic Structure Calculation. Part B. Applications, volume 8 of Modern theoretical chemistry. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-33508-5. ISSN 0146-0420. xv + 308 pp. LCCN QD461 .S396.

Slater:1933:ITP

[SF33] John C. (John Clarke) Slater and Nathaniel Herman Frank. Introduction to Theoretical Physics. International series in physics. McGraw-Hill, New York, NY, USA, 1933. xx + 576 pp. LCCN QC21 .S595 1933c.

Slater:1947:E

[SF47a] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1947. xiii + 240 pp. LCCN QC760 .S55.

Slater:1947:M

[SF47b] John C. (John Clarke) Slater and Nathaniel Herman Frank. Mechanics. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1947. xiii + 297 pp. LCCN QA805 .S632.

Slater:1951:IFS

[SF51] John C. (John Clarke) Slater and Nathaniel Herman Frank. Introducción a La Física Teórica. (Spanish) [Introduction to Theoretical Physics]. Espasa-Calpe Argentina, Buenos Aires, Argentina, 1951. ???? pp. Translation to Spanish by Rafael Grinfeld (1902–1969).

Slater:1969:E

[SF69] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. Dover, New York, NY, USA, 1969. ISBN 0-486-62263-0. xiii + 240 pp. LCCN QC760 .S55 1969. URL http://

www.loc.gov/catdir/description/dover032/69017476.html.

Slater:1973:E

[SF73] John C. (John Clarke) Slater and Nathaniel Herman Frank. Electromagnetism. Haryana Hindi Granth Akademi, Chandigarh, Punjab, India, 1973. xvi + 256 pp.

Slater:1983:M

[SF83] John C. (John Clarke) Slater and Nathaniel Herman Frank. Mechanics. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1983. ISBN 0-313-24064-7 (hardcover). xiii + 297 pp. LCCN QA805 .S632 1983.

Stewart:2013:SAU

[SHR13] Brianna Stewart, Derrick J. Hylton, and Natarajan Ravi. A systematic approach for understanding Slater–Gaussian functions in computational chemistry. Journal of Chemical Education, 90(5): 609–612, April 2013. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Slater:1972:SCF

[SJ72] J. C. Slater and K. H. Johnson. Self-consistent-field X α cluster method for polyatomic molecules and solids. Physical Review B: Solid State, 5(3):844–853, February 1, 1972. CODEN PLRBAQ. ISSN 0556-2805.

Slater:1974:QCC

[SJ74] John C. Slater and Keith H. Johnson. Quantum chemistry and catalysis. Physics Today, 27(10):34–41, October 1974. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v27/
i10/p34_s1.

Slater:1931:VWF

[SK31] John C. Slater and John G. Kirkwood. The van der Waals forces in gases. Physical Review (2), 37(6):682–697, March 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.37.
682;   http://prola.aps.org/abstract/PR/v37/i6/p682_1.

Slater:1935:TFM

[SK35] J. C. Slater and H. M. Krutter. The Thomas–Fermi method for metals. Physical Review (2), 47(7):559–568, April 1, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v47/i7/p559_1.

Slater:1954:SLM

[SK54] J. C. Slater and G. F. Koster. Simplified LCAO method for the periodic potential problem. Physical Review (2), 94(6):1498– 1524, June 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v94/i6/p1498_1.

Slater:1962:SFE

[SKW62] J. C. Slater, G. F. Koster, and J. H. Wood. Symmetry and free electron properties of the gallium energy bands. Physical Review (2), 126(4):1307–1317, May 15, 1962. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v126/i4/p1307_1.

Slater:1920:DPI

[Sla20] John Clarke Slater. Dependence of pressure of the intensities of the Balmer lines of hydrogen. Honors B.S. thesis, Harvard University, Cambridge, MA, USA, 1920. Title and date uncertain: not in Harvard Hollis Library catalog, but cited in memoir [Mor82].

Slater:1923:CAH

[Sla23] John Clarke Slater. The Compressibility of the Alkali-Halides. Ph.D. thesis, Harvard University, Cambridge, MA, USA, 1923.

URL http://search.proquest.com/docview/301815314.

Slater:1924:CAH

[Sla24a] J. C. Slater. Compressibility of the alkali halides. Physical Review (2), 23(4):488–500, April 1, 1924. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v23/i4/p488_1.

Slater:1924:RA

[Sla24b] J. C. Slater. Radiation and atoms. Nature, 113(2835):307–308, March 1, 1924. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v113/n2835/pdf/113307b0.pdf.

Slater:1925:IHH

[Sla25a] J. C. Slater. Interpretation of the hydrogen and helium spectra. Proceedings of the National Academy of Sciences of the United States of America, 11(12):732–738, December 15, 1925. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL

http://www.jstor.org/stable/84697.

Slater:1925:MDT

[Sla25b] J. C. Slater. Methods for determining transition probabilities from line absorption. Physical Review (2), 25(6):783–790, June 1, 1925. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v25/i6/
p783_1.

Slater:1925:NR

[Sla25c] J. C. Slater. The nature of radiation. Nature, 116(2912):278, August 22, 1925. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v116/n2912/pdf/116278a0.pdf.

Slater:1925:PDS

[Sla25d] J. C. Slater. Physically degenerate systems and quantum dynamics. Physical Review (2), 26(4):419–430, October 1, 1925. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v26/i4/p419_1.

Slater:1925:QTO

[Sla25e] J. C. Slater. A quantum theory of optical phenomena. Physical Review (2), 25(4):395–428, April 1, 1925. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v25/i4/p395_1.

Slater:1926:AIB

[Sla26a] J. C. Slater. Alternating intensities in band lines. Nature, 117 (2946):555–556, April 17, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v117/n2946/pdf/117555b0.pdf.

Slater:1926:DMC

[Sla26b] J. C. Slater. A dynamical model for complex atoms. Physical Review (2), 28(2):291–317, August 1, 1926. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v28/i2/p291_1.

Slater:1926:MCA

[Sla26c] J. C. Slater. Measurement of the compressibility of the alkali halides. Proceedings of the American Academy of Arts and Sciences, 61(5):135–150, April 1926. CODEN PAAAAV. ISSN 0065-

6836. URL http://www.jstor.org/stable/20026146.

Slater:1926:SES

[Sla26d] J. C. Slater. Spinning electrons and the structure of spectra. Nature, 117(2947):587, April 24, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v117/n2947/pdf/117587a0.pdf.

Slater:1927:ARP

[Sla27a] J. C. Slater. Action of radiation and perturbations on atoms. Proceedings of the National Academy of Sciences of the United States of America, 13(3):104–111, March 15, 1927. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

jstor.org/stable/85161.

Slater:1927:RAS

[Sla27b] J. C. Slater. Radiation and absorption on Schrödinger’s theory. Proceedings of the National Academy of Sciences of the United States of America, 13(1):7–12, January 15, 1927. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL

http://www.jstor.org/stable/85066.

Slater:1927:SHA

[Sla27c] J. C. Slater. The structure of the helium atom. I. Proceedings of the National Academy of Sciences of the United States of America, 13(6):423–430, June 15, 1927. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/84939.

Slater:1928:CFR

[Sla28a] J. C. Slater. Central fields and Rydberg formulas in wave mechanics. Physical Review (2), 31(3):333–343, March 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v31/i3/p333_1.

Slater:1928:LQW

[Sla28b] J. C. Slater. Light quanta and wave mechanics. Physical Review (2), 31(5):895–899, May 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v31/i5/p895_1.

Slater:1928:NSH

[Sla28c] J. C. Slater. The normal state of helium. Physical Review (2), 32 (3):349–360, September 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v32/i3/p349_1.

Slater:1928:SCF

[Sla28d] J. C. Slater. The self consistent field and the structure of atoms. Physical Review (2), 32(3):339–348, September 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v32/i3/p339_1.

Slater:1929:TCS

[Sla29a] J. C. Slater. The theory of complex spectra. Physical Review (2), 34(10):1293–1322, November 15, 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v34/i10/p1293_1.

Slater:1929:PMW

[Sla29b] John C. Slater. Physical meaning of wave mechanics. Journal of The Franklin Institute, 207(4):449–455, April 1929. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0016003229918238.

Slater:1930:ASC

[Sla30a] J. C. Slater. Atomic shielding constants. Physical Review (2), 36(1):57–64, July 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v36/i1/p57_1.

Slater:1930:CMM

[Sla30b] J. C. Slater. Cohesion in monovalent metals. Physical Review (2), 35(5):509–529, March 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v35/i5/p509_1. See comment [Löw66a].

Reprinted in [Löw66b, pages 17–??].

Slater:1930:NHM

[Sla30c] J. C. Slater. Note on Hartree’s method. Physical Review (2), 35 (2):210–211, January 15, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v35/i2/p210_2.

Slater:1931:DVP

[Sla31a] J. C. Slater. Directed valence in polyatomic molecules. Physical Review (2), 37(5):481–489, March 1, 1931. CODEN PHRVAO.

ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v37/i5/p481_1. See also [Pau31a,

Pau31b].

Slater:1931:MEL

[Sla31b] J. C. Slater. Molecular energy levels and valence bonds. Physical Review (2), 38(6):1109–1144, September 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v38/i6/p1109_1.

Slater:1931:NSG

[Sla31c] J. C. Slater. Note on the structure of the groups X O_{3}. Physical Review (2), 38(2):325–329, July 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v38/i2/p325_1.

Slater:1931:QTE

[Sla31d] J. C. Slater. The quantum theory of the equation of state. Physical Review (2), 38(2):237–242, July 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v38/i2/p237_1.

Slater:1931:QTM

[Sla31e] John Clarke Slater. Quantum Theory of Matter. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1931. 528 pp. LCCN QC174.1 S63.

Slater:1932:AAW

[Sla32a] J. C. Slater. Analytic atomic wave functions. Physical Review (2), 42(1):33–43, October 1, 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v42/i1/p33_1.

Slater:1932:NMS

[Sla32b] John C. Slater. Note on molecular structure. Physical Review (2), 41(2):255–257, July 15, 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.41.255.2;   http://
prola.aps.org/abstract/PR/v41/i2/p255_2.

Slater:1933:ETM

[Sla33a] J. C. Slater. The electron theory of metallic conduction. Science, 77 (2008):595–597, June 23, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1658314;   http://www.sciencemag.org/content/77/
2008/595.full.pdf.

Slater:1933:VMS

[Sla33b] John C. Slater. The virial and molecular structure. Journal of Chemical Physics, 1(10):687–691, October 1, 1933. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://jcp.aip.org/resource/1/jcpsa6/v1/i10/p687_s1.

Slater:1934:EEB

[Sla34a] J. C. Slater. Electronic energy bands in metals. Physical Review (2), 45(11):794–801, June 1, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v45/i11/p794_1.

Slater:1934:ESM

[Sla34b] J. C. Slater. The electronic structure of metals. Reviews of Modern Physics, 6(4):209–280, October 1, 1934. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.6.
209;   http://rmp.aps.org/abstract/RMP/v6/i4/p209_1.

Slater:1936:FN

[Sla36a] J. C. Slater. The ferromagnetism of nickel. Physical Review (2), 49(7):537–545, April 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v49/i7/p537_1.

Slater:1936:FNI

[Sla36b] J. C. Slater. The ferromagnetism of nickel. II. temperature effects. Physical Review (2), 49(12):931–937, June 15, 1936. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v49/i12/p931_1.

Slater:1937:DEW

[Sla37a] J. C. Slater. Damped electron waves in crystals. Physical Review (2), 51(10):840–846, May 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v51/i10/p840_1.

Slater:1937:NSSa

[Sla37b] J. C. Slater. The nature of the superconducting state. Physical Review (2), 51(3):195–202, February 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v51/i3/p195_1.

Slater:1937:NSSb

[Sla37c] J. C. Slater. The nature of the superconducting state. II. Physical Review (2), 52(3):214–222, August 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v52/i3/p214_1.

Slater:1937:TFL

[Sla37d] J. C. Slater. The theory of ferromagnetism: Lowest energy levels. Physical Review (2), 52(3):198–214, August 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v52/i3/p198_1.

Slater:1937:WFP

[Sla37e] J. C. Slater. Wave functions in a periodic potential. Physical Review (2), 51(10):846–851, May 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v51/i10/p846_1.

Slater:1937:CEM

[Sla37f] John C. Slater. Charles Elwood Mendenhall (1872–1935). Proceedings of the American Academy of Arts and Sciences, 71(10): 529–530, March 1937. CODEN PAAAAV. ISSN 0065-6836. URL

http://www.jstor.org/stable/20023261.

Slater:1937:ESA

[Sla37g] John C. Slater. Electronic structure of alloys. Journal of Applied Physics, 8(6):385–390, June 1, 1937. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http:/

/jap.aip.org/resource/1/japiau/v8/i6/p385_s1.

Slater:1937:TCO

[Sla37h] John C. Slater. Thermodynamical considerations of order and disorder. In Anonymous [Ano37], page ?? LCCN ????

Slater:1938:EPB

[Sla38a] John C. Slater. Electrodynamics of ponderable bodies. Journal of The Franklin Institute, 225(3):277–287, March 1938. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0016003238907021.

Slater:1938:EEL

[Sla38b] John C. Slater. Excited energy levels of insulating crystals. Transactions of the Faraday Society, 34(??):828–832, ???? 1938. CO-DEN TFSOA4. ISSN 0014-7672. URL http://pubs.rsc.org/

en/content/articlelanding/1938/tf/tf9383400828.

Slater:1939:ICP

[Sla39] John C. (John Clarke) Slater. Introduction to Chemical Physics. International series in physics; F. K. Richtmeyer, consulting editor. McGraw-Hill, New York, NY, USA, 1939. xiv + 521 pp. LCCN QC171 .S55.

Slater:1940:NGC

[Sla40a] J. C. Slater. Note on Grüneisen’s constant for the incompressible metals. Physical Review (2), 57(8):744–746, April 15, 1940. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v57/i8/p744_1.

Slater:1940:NEP

[Sla40b] J. C. Slater. Note on the effect of pressure on the Curie point of iron–nickel alloys. Physical Review (2), 58(1):54–56, July 1, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v58/i1/
p54_1.

Slater:1941:TTK

[Sla41] John C. Slater. Theory of the transition in KH_{2}PO_{4}. Journal of Chemical Physics, 9(1):16–33, January 1, 1941. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http:/

/jcp.aip.org/resource/1/jcpsa6/v9/i1/p16_s1. Republished

in [Sla87].

Slater:1942:MT

[Sla42] John C. (John Clarke) Slater. Microwave Transmission. International series in physics. McGraw-Hill, New York, NY, USA, 1942. x + 309 pp. LCCN QC661 .S635.

Slater:1946:ME

[Sla46a] J. C. Slater. Microwave electronics. Reviews of Modern Physics, 18(4):441–512, October 1, 1946. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.18.441;   http://
rmp.aps.org/abstract/RMP/v18/i4/p441_1.

Slater:1946:PWE

[Sla46b] J. C. Slater. Physics and the wave equation. Bulletin of the American Mathematical Society, 52(??):392–400, 1946. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). The nineteenth Josiah Willard Gibbs lecture delivered at Chicago, Illinois, November 23, 1945.

Slater:1947:PUR

[Sla47] John C. (John Clarke) Slater. Peredacha ulstrakorotkikh radiovola. ????, ????, 1947. 344 pp. LCCN QC661 .S639 1947.

Slater:1948:DLA

[Sla48a] J. C. Slater. The design of linear accelerators. Reviews of Modern Physics, 20(3):473–518, July 1, 1948. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.20.473;
 http://rmp.aps.org/abstract/RMP/v20/i3/p473_1.

Slater:1948:HMP

[Sla48b] John C. Slater. History of the M.I.T. Physics Department, 1930– 48. Manuscript in Slater Papers, American Philosophical Society Library, Philadelphia, PA., 1948.

Slater:1949:EPP

[Sla49a] J. C. Slater. Electrons in perturbed periodic lattices. Physical Review (2), 76(11):1592–1601, December 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v76/i11/p1592_1.

Slater:1949:IAC

[Sla49b] John C. Slater. Introduzione Alla Chimica Fisica. (Italian) [Introduction to Physical Chemistry]. Sansoni Edizioni Scientifiche, Firenze, Italia, 1949. x + 560 pp. LCCN 541 SL1I.I. Translation to Italian by Eolo Scrocco.

Slater:1949:PM

[Sla49c] John C. Slater. The physics of metals. Physics Today, 2(1):6– 13, January 1949. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v2/i1/p6_s1. See letter [Sla49d].

Slater:1949:TA

[Sla49d] John C. Slater. Thanks again. Physics Today, 2(3):3, March 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v2/i3/p3_s1.

Slater:1949:ICP

[Sla49e] John C. Slater. The MIT International Conference on the Physics of Very Low Temperatures, September 6–10, 1949. Science, 110(2862):465–467, November 4, 1949. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/www.jstor.org/stable/1676026;   http://www.sciencemag.
org/content/110/2862/465.full.pdf.

Slater:1950:LCB

[Sla50a] J. C. Slater. The Lorentz correction in barium titanate. Physical Review (2), 78(6):748–761, June 15, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v78/i6/p748_1.

Slater:1950:SPA

[Sla50b] John C. Slater. Structure and polarization of atoms and molecules. Electrical Engineering (American Institute of Electrical Engineers), 69(??):855–858, ???? 1950. CODEN ELENAC. ISSN 0095-9197.

Slater:1950:ME

[Sla50c] John C. (John Clarke) Slater. Microwave Electronics. The Bell Telephone Laboratories series. Van Nostrand, New York, NY, USA, 1950. xiv + 406 pp. LCCN QC721 .S6.

Slater:1951:ECC

[Sla51a] J. C. Slater. The effect of chemical combination on the internal conversion in Tc. Physical Review (2), 84(6):1261–1262, December 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v84/i6/
p1261_1.

Slater:1951:ETS

[Sla51b] J. C. Slater. The electron theory of solids. American Journal of Physics, 19(6):368–374, September 1951. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.

org/resource/1/ajpias/v19/i6/p368_s1.

Slater:1951:MEH

[Sla51c] J. C. Slater. Magnetic effects and the Hartree–Fock equation. Physical Review (2), 82(4):538–541, May 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v82/i4/p538_1.

Slater:1951:NSB

[Sla51d] J. C. Slater. Note on superlattices and Brillouin zones. Physical Review (2), 84(2):179–181, October 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v84/i2/p179_1.

Slater:1951:SHF

[Sla51e] J. C. Slater. A simplification of the Hartree–Fock method. Physical Review (2), 81(3):385–390, February 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v81/i3/p385_1.

Slater:1951:DWF

[Sla51f] John C. Slater. Determinantal wave functions and magnetic problems. In ????, editor, Conference on Quantum-Mechanical Methods in Valence Theory, pages 121–126. Office of Naval Research, Washington, DC, USA, 1951. LCCN ????

Slater:1951:ERM

[Sla51g] John C. Slater. The effects of radiation on materials. Journal of Applied Physics, 22(3):237–256, March 1, 1951. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL

http://link.aip.org/link/japiau/v22/i3/p237/s1.

Slater:1951:NOA

[Sla51h] John C. Slater. Note on orthogonal atomic orbitals. Journal of Chemical Physics, 19(2):220–223, February 1, 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://jcp.aip.org/resource/1/jcpsa6/v19/i2/p220_s1.

Slater:1951:RSC

[Sla51i] John C. Slater. Remarks on self-consistent molecular orbitals. In ????, editor, Conference on Quantum-Mechanical Methods in Valence Theory, page 185. Office of Naval Research, Washington, DC, USA, 1951. LCCN ????

Slater:1951:QTM

[Sla51j] John Clarke Slater. Quantum Theory of Matter. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1951. xiv + 528 pp. LCCN QC174.1 S63.

Slater:1952:SPE

[Sla52a] J. C. Slater. A soluble problem in energy bands. Physical Review (2), 87(5):807–835, September 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v87/i5/p807_1.

Slater:1952:HEA

[Sla52b] John C. Slater. High energy accelerators: Linear accelerators. Annual Review of Astronomy and Astrophysics, 1:199–206, 1952. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-4282

(electronic). URL http://www.annualreviews.org/doi/abs/10.
1146/annurev.ns.01.120152.001215.

Slater:1952:IEC

[Sla52c] John C. Slater. IUPAP Executive Committee. Physics Today, 5 (9):32–33, September 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v5/i9/p32_s1?isAuthorized=no.

Slater:1952:SS

[Sla52d] John C. Slater. The solid state. Physics Today, 5(1):10–15, January 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v5/i1/p10_s1.

Slater:1953:APW

[Sla53a] J. C. Slater. An augmented plane wave method for the periodic potential problem. Physical Review (2), 92(3):603–608, November 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v92/i3/
p603_1.

Slater:1953:FBT

[Sla53b] J. C. Slater. Ferromagnetism and the band theory. Reviews of Modern Physics, 25(1):199–210, January 1, 1953. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.25.199;
 http://rmp.aps.org/abstract/RMP/v25/i1/p199_1.

Slater:1953:GSC

[Sla53c] J. C. Slater. A generalized self-consistent field method. Physical Review (2), 91(3):528–530, August 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v91/i3/p528_1.

Slater:1953:ESA

[Sla53d] John C. Slater. Electronic structure of atoms and molecules. Report, Solid State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, February 1953. iv + 212 pp.

Slater:1953:ESSa

[Sla53e] John C. Slater. Electronic structure of solids. I. The energy band method. Report 4, Solid State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, July 1953.

Slater:1953:ESSb

[Sla53f] John C. Slater. Electronic structure of solids. II. The perturbed periodic lattice. Report 5, Solid State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, December 1953.

Slater:1953:PF

[Sla53g] John C. Slater. Problem of ferromagnetism. In ????, editor, Proceedings of the International Conference on Theoretical Physics, Kyoto and Tokyo, pages 679–693. ????, ????, 1953. LCCN ????

Slater:1953:WMTa

[Sla53h] John C. Slater. Work on molecular theory in the Solid State and Molecular Theory Group, MIT. In ????, editor, Symposium on Molecular Physics, Nikko, pages 1–4. ????, ????, 1953. LCCN ????

Slater:1953:WMTb

[Sla53i] John C. Slater. Work on molecular theory in the Solid State and Molecular Theory Group, MIT. In ????, editor, Proceedings of the International Conference on Theoretical Physics, Kyoto and Tokyo, pages 611–621. ????, ????, 1953. LCCN ????

Slater:1953:QTM

[Sla53j] John Clarke Slater. Quantum Theory of Matter. McGraw-Hill, New York, NY, USA, 1953. xiv + 528 pp.

Slater:1954:ESS

[Sla54] John C. Slater. Electronic structure of solids. III. Configuration interaction in solids. Report 6, Solid State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, April 1954.

Slater:1955:OEE

[Sla55a] J. C. Slater. One-electron energies of atoms, molecules, and solids. Physical Review (2), 98(4):1039–1045, May 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v98/i4/p1039_1.

Slater:1955:MP

[Sla55b] John C. (John Clarke) Slater. Modern Physics. McGraw-Hill, New York, NY, USA, 1955. xi + 322 pp. LCCN QC7 .S57.

Slater:1956:BTP

[Sla56a] J. C. Slater. Barrier theory of the photoconductivity of lead sulfide. Physical Review (2), 103(6):1631–1644, September 15, 1956. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v103/i6/p1631_1.

Slater:1956:EAM

[Sla56b] J. C. Slater. Book review: Electrons, atoms, metals, and alloys, William Hume-Rothery, Royal Society Warren Research Fellow, Lecturer in Metallurgical Chemistry, University of Oxford. Revised edition. The Philosophical Library, New York. 1955. 387pp. 171 figs. 10 tables. 14.5 × 22 cm. $10. Journal of Chemical Education, 33(10):A470, A472, October 1956. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Slater:1956:BTB

[Sla56c] John C. Slater. Band theory of bonding in metals. In ????, editor, Theory of Alloy Phases: a seminar on theory of alloy phases held during the Thirty-Seventh National Metal Congress and Exposition, Philadelphia, October 15 to 21, 1955, pages 1–12. American Society for Metals, Cleveland, OH, USA, 1956. LCCN QD135 A51 1955.

Slater:1956:BRB

[Sla56d] John C. Slater. Book review: Electrons, atoms, metals, and alloys, by William Hume-Rothery. Journal of Chemical Education, 33(10): A470, A472, October 1956. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL http://pubs.acs.org/doi/

abs/10.1021/ed033pA470.

Slater:1956:MRP

[Sla56e] John C. Slater. Materials research program. Slater papers, folder MIT Materials Research #1., 1956.

Slater:1956:PEP

[Sla56f] John C. Slater. Proposal for an expanded program of materials research at the Massachusetts Institute of Technology, July 12, 1956. Slater papers, folder MIT Materials Research #1., July 12, 1956.

Slater:1956:ESS

[Sla57] John C. (John Clarke) Slater. The electronic structure of solids. In Siegfried Flügge., editor, Handbuch der Physik: Elektrische Leitungsphänomene. (German) [Electrical conductivity phenomena], volume 19–20, pages 1–136. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1956/1957. ISBN 0-387-02040-3, 3-540-02040-3. LCCN QC21 .H327.

Slater:1958:ICC

[Sla58a] J. C. Slater. International conference on current problems in crystal physics. Reviews of Modern Physics, 30(1):46, January 1, 1958. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.30.46;   http://rmp.aps.org/
abstract/RMP/v30/i1/p46_1.

Slater:1958:IWC

[Sla58b] John C. Slater. Interaction of waves in crystals. Reviews of Modern Physics, 30(1):197–222, January 1, 1958. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.30.197;
 http://rmp.aps.org/abstract/RMP/v30/i1/p197_1.

Slater:1959:BT

[Sla59a] John C. Slater. Band theory. The Journal of Physics and Chemistry of Solids, 8:21–25, January 1959. CODEN JPCSAW. ISSN 0022-3697 (print), 1879-2553 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022369759902653.

Slater:1959:MT

[Sla59b] John C. Slater. Microwave Transmission. Dover, New York, NY, USA, 1959. x + 309 pp.

Slater:1959:NIS

[Sla59c] John C. Slater. Note on the interatomic spacings in the ions I_{3}^{−}, FHF^{−}. Acta Crystallographica, 12(3):197–200, March 1959. CO-DEN ACCRA9. ISSN 0365-110X (print), 1600-8642 (electronic).

Slater:1960:MT

[Sla60a] J. C. Slater. Microwave Transmission. Dover, New York, NY, USA, 1960. x + 309 pp.

Slater:1960:MC

[Sla60b] John C. Slater. On the MIT Materials Center. Slater papers, folder MIT Materials Research #1., 1960.

Slater:1960:QTA

[Sla60c] John C. (John Clarke) Slater. Quantum Theory of Atomic Structure. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1960. ISBN 0-07-058040-5. xii + 502 (volume 1), ix + 439 (volume 2) pp. LCCN QC174.1 .S53.

Slater:1962:AM

[Sla62] John C. Slater. Administrative memo #1. Slater papers, folder MIT Materials Research #1., July 20, 1962.

Slater:1963:DAS

[Sla63a] J. C. Slater. Discussion on atomic spectra. Reviews of Modern Physics, 35(3):520–521, July 1, 1963. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.35.520;
 http://rmp.aps.org/abstract/RMP/v35/i3/p520_1.

Slater:1963:ESA

[Sla63b] J. C. Slater. The electronic structure of atoms — the Hartree– Fock method and correlation. Reviews of Modern Physics, 35 (3):484–487, July 1, 1963. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.35.484;   http://
rmp.aps.org/abstract/RMP/v35/i3/p484_1.

Slater:1963:QTM

[Sla63c] John C. Slater. Quantum Theory of Molecules and Solids: Electronic-Structure of Molecules, volume 1. McGraw-Hill, New York, NY, USA, 1963. xv + 485 pp. LCCN QC174.1 .S551.

Slater:1964:BRC

[Sla64a] J. C. Slater. Book review: C. Kittel, Quantum Theory of Solids. American Scientist, 52(4):451A–452A, December 1964. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27839196.

Slater:1964:EBC

[Sla64b] J. C. Slater. Energy band calculations by the Augmented Plane Wave method. Advances in Quantum Chemistry, 1:35–58, 1964. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327608603743.

Slater:1964:ARC

[Sla64c] John C. Slater. Atomic radii in crystals. Journal of Chemical Physics, 41(10):3199–3204, November 15, 1964. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://jcp.

aip.org/resource/1/jcpsa6/v41/i10/p3199_s1.

Slater:1964:RMN

[Sla64d] John C. Slater. Robert Mulliken of Newburyport. In Löwdin and Pullman [LP64], pages 17–20. LCCN QD461 .L872 1964.

Slater:1965:QTM

[Sla65a] J. C. Slater. Quantum Theory of Molecules and Solids. Volume 2: Symmetry and Energy Bands in Crystals. McGraw-Hill, New York, NY, USA, 1965. xii + 563 pp. LCCN ????

Slater:1965:SGW

[Sla65b] J. C. Slater. Space groups and wave-function symmetry in crystals. Reviews of Modern Physics, 37(1):68–83, January 1, 1965. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
37.68;   http://rmp.aps.org/abstract/RMP/v37/i1/p68_1.

Slater:1965:MOH

[Sla65c] John C. Slater. Molecular orbital and Heitler–London methods (with discussion). Journal of Chemical Physics, 43(10):S11–S17, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v43/i10/pS11_s1. See discussion [Har65].

Slater:1965:PWB

[Sla65d] John C. Slater. P. W. Bridgman and high-pressure physics. Science, 148(3671):805–806, May 7, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/www.jstor.org/stable/1716369;   http://www.sciencemag.
org/content/148/3671/805.full.pdf.

Slater:1965:SSS

[Sla65e] John C. Slater. Suggestions from solid state theory regarding molecular calculations. Journal of Chemical Physics, 43(10):S228, November 15, 1965. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v43/i10/pS228_s1.

Slater:1966:GFM

[Sla66] J. C. Slater. Green’s-function method in the energy-band problem. Physical Review (2), 145(2):599–602, May 13, 1966. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v145/i2/p599_1.

Slater:1967:CPS

[Sla67a] J. C. Slater. Correlation problems in solids. International Journal of Quantum Chemistry, 1(S1):783–789, January 16–21, 1967. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory.

Slater:1967:EBF

[Sla67b] J. C. Slater. Energy bands and Fermi surfaces. International Journal of Quantum Chemistry, 1(S1):523–531, January 16–21, 1967. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory.

Slater:1967:QPA

[Sla67c] J. C. Slater. Quantum physics in America between the wars. International Journal of Quantum Chemistry, 1(S1):1–23, January 16–21, 1967. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory.

Slater:1967:BRA

[Sla67d] John C. Slater. Book review: APW calculations for the beginner: Terry Loucks, Augmented Plane Wave Method: A Guide to Performing Electronic Structure Calculations. Physics Today, 20 (11):92–93, November 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v20/i11/p92_s1.

Slater:1967:CSS

[Sla67e] John C. Slater. The current state of solid-state and molecular theory. International Journal of Quantum Chemistry, 1(1):37–102, January 1967. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1967:ITF

[Sla67f] John C. Slater. Introduction to the theory of ferroelectricity. In Weller [Wel67], pages 1–8. LCCN QC595 S98.

Slater:1967:QTM

[Sla67g] John Clarke Slater. Quantum Theory of Molecules and Solids: Insulators, Semiconducters, and Metals, volume 3 of International series in pure and applied physics, 99-0107607-2. McGraw-Hill, New York, NY, USA, 1967. xiv + 549 pp. LCCN QC 174.12.

Slater:1968:AES

[Sla68a] J. C. Slater. Average energy of states of given multiplicities in atoms. Physical Review (2), 165(2):655–658, January 10, 1968. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v165/i2/p655_1.

Slater:1968:BRS

[Sla68b] J. C. Slater. Book review: Second Quantization and Atomic Spectroscopy, Brian R. Judd, pp. 61 + viii, the Johns Hopkins Press, Baltimore, 1967. Price $5.95. American Journal of Physics, 36 (1):69, January 1968. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL https://aapt.scitation.org/doi/
10.1119/1.1974426.

Slater:1968:ESP

[Sla68c] J. C. Slater. Exchange in spin-polarized energy bands. Physical Review (2), 165(2):658–669, January 10, 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v165/i2/p658_1.

Slater:1968:EBTa

[Sla68d] John C. Slater. Energy band theory of magnetism. Journal of Applied Physics, 39(2):761–767, February 1, 1968. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL

http://link.aip.org/link/japiau/v39/i2/p761/s1.

Slater:1968:EBTb

[Sla68e] John C. Slater. Energy bands and the theory of solids. In Alder et al. [AFR68], pages 1–20. LCCN QA401 .M514 v.8.

Slater:1968:EBS

[Sla68f] John C. Slater. Energy bands in solids. Physics Today, 21(4):61–71, April 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v21/i4/p61_s1.

Slater:1968:QPA

[Sla68g] John C. Slater. Quantum physics in America between the wars. Physics Today, 21(1):43–51, January 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v21/i1/p43_s1.

Slater:1968:QTM

[Sla68h] John C. Slater. Quantum Theory of Matter. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, second edition, 1968. x + 763 pp. LCCN QC174.1 S63.

Slater:1968:REB

[Sla68i] John C. Slater. Review of the energy band problem with recent results. In L. H. (Lawrence Herman) Bennett and J. T. (James Thomas) Waber, editors, Energy Bands in Metals and Alloys, volume 45 of Metallurgical Society conferences, pages 1–17. Gordon and Breach, New York, NY, USA, 1968. LCCN ???? Papers of a symposium sponsored by the I.M.D. Committee on Alloy Phases of the Metallurgical Society, held February 13, 1967, at the ninety-sixth annual meeting of the AIME in Los Angeles.

Slater:1969:SCF

[Sla69a] J. C. Slater. The self-consistent field for crystals. International Journal of Quantum Chemistry, 4(S3B):727–746, January 13–18, 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Slater:1969:WC

[Sla69b] John C. Slater. Why has he changed? Physics Today, 22 (3):35–37, March 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v22/i3/p35_s1.

Slater:1969:CDQ

[Sla69c] John C. (John Clarke) Slater. Concepts and Development of Quantum Physics. Dover, New York, NY, USA, 1969. ISBN 0-486-62265-7. xi + 322 pp. LCCN QC174.1 .S528 1969. Reprint of [Sla55b].

Slater:1969:ME

[Sla69d] John C. (John Clarke) Slater. Microwave Electronics. Dover, New York, NY, USA, 1969. ISBN 0-486-62264-9. xiv + 406 pp. LCCN TK7876 .S58 1969.

Slater:1970:NSP

[Sla70a] John C. Slater. Note on the space part of anti-symmetric wave functions in the many-electron problem. International Journal of Quantum Chemistry, 4(6):561–570, November 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1970:PSX

[Sla70b] John C. Slater. Present status of the Xα statistical exchange. Report 71, Solid State and Molecular Theory Group, Massachusetts Institute of Technology, Cambridge, MA, USA, July 1970.

Slater:1970:ICP

[Sla70c] John C. (John Clarke) Slater. Introduction to Chemical Physics. Dover, New York, NY, USA, 1970. ISBN 0-486-62562-1. xiv + 521 pp. LCCN QC171 .S55 1970.

Slater:1971:TEA

[Sla71a] J. C. Slater. Treatment of exchange in atomic, molecular, and solidstate theory. International Journal of Quantum Chemistry, 5(S5): 403–409, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1971:IPJ

[Sla71b] John C. Slater. Introduction of Professor J. H. Van Vleck, Sanibel Symposium, Monday, January 18, 1971. International Journal of Quantum Chemistry, 5(S5):1–2, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1971:SCF

[Sla71c] John C. Slater. The self-consistent field method for crystals. In Marcus et al. [MJW71], pages 447–457. ISBN 1-4684-1892-0 (print), 1-4684-1890-4 (e-book). URL http://link.springer.

com/book/10.1007/978-1-4684-1890-3.

Slater:1971:TPF

[Sla71d] John C. Slater. Transition probabilities and fractional occupation numbers in atoms. In Brittin and Odabasi [BO71], pages 297–303. ISBN 0-85274-195-2, 0-87081-010-3. LCCN QC71 .T63.

Slater:1972:HFV

[Sla72a] John C. Slater. Hellmann–Feynman and virial theorems in the X α method. Journal of Chemical Physics, 57(6):2389–2396, September 15, 1972. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://jcp.aip.org/resource/1/
jcpsa6/v57/i6/p2389_s1.

Slater:1972:NMT

[Sla72b] John C. Slater. New methods in the theory of molecules and solids. Journal de physique, 33(C3):1–5, May/June 1972. CODEN JOPQAG. ISSN 0302-0738.

Slater:1972:SEH

[Sla72c] John C. Slater. Statistical exchange and the Heisenberg exchange integral. Journal de physique, 33(C3):7–11, ???? 1972. CODEN JOPQAG. ISSN 0302-0738.

Slater:1972:SEC

[Sla72d] John C. Slater. Statistical exchange-correlation in the selfconsistent field. Advances in Quantum Chemistry, 6:1–92, 1972. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327608605419.

Slater:1972:SEB

[Sla72e] John C. (John Clarke) Slater. Symmetry and Energy Bands in Crystals. Dover, New York, NY, USA, 1972. ISBN 0-486-60573-6. xii + 563 pp. LCCN QD911 .S57 1972.

Slater:1973:FPX

[Sla73a] J. C. Slater. Future prospects for the X_{α} method. International Journal of Quantum Chemistry, 7(S7):533–544, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1973:IUC

[Sla73b] J. C. Slater. Introduction of E. U. Condon, Sanibel Symposium January 22, 1973. International Journal of Quantum Chemistry, 7 (S7):3–5, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Slater:1973:DQM

[Sla73c] John C. Slater. The development of quantum mechanics in the period 1924–1926. In Price et al. [PCR73], pages 19–25. ISBN 0-470-69731-8. LCCN QC174.22 .W38 1973.

Slater:1974:RIM

[Sla74a] J. C. Slater. Recent improvements in the muffin-tin method. International Journal of Quantum Chemistry, 8(S8):81–88, January 20–26, 1974. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Slater:1974:HXM

[Sla74b] John C. Slater. The history of the Xα method. In Daudel and Pullman [DP74], page ?? ISBN 90-277-0421-X. LCCN QD462.A1 I57 1973.

Slater:1974:QTM

[Sla74c] John C. (John Clarke) Slater. Quantum Theory of Molecules and Solids: The Self-Consistent Field for Molecules and Solids. International series in pure and applied physics. McGraw-Hill, New York, NY, USA, 1974. ISBN 0-07-058038-3. xvii + 583 pp. LCCN QC174.12 .S55.

Slater:1975:CTX

[Sla75a] J. C. Slater. Comparison of TFD and Xα methods for molecules and solids. International Journal of Quantum Chemistry, 9(S9): 7–21, January 19–25, 1975. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Statistics.

Slater:1975:BRF

[Sla75b] John C. Slater. Book review: F. Hund, The History of Quantum Theory. Physics Today, 28(8):69, August 1975. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v28/i8/p69_s1.

Slater:1975:TPL

[Sla75c] John C. Slater. Typescript of “A Physicist of the Lucky Generation”. The text is not available at the MIT Web site., 1975. URL

https://archivesspace.mit.edu/repositories/2/resources/
723;   https://snaccooperative.org/vocab_administrator/
resources/6358261.

Slater:1975:SSM

[Sla75d] John C. (John Clarke) Slater. Solid-state and Molecular Theory: a Scientific Biography. Wiley, New York, NY, USA, 1975. ISBN 0-471-79681-6. viii + 357 pp. LCCN QC176.8.E4 S55 1975.

Slater:1976:ILB

[Sla76a] John C. Slater. Interdisciplinary laboratories for basic research in materials sciences. Slater papers, folder MIT Dept. of Physics #39., 1976.

Slater:1976:PSM

[Sla76b] John C. Slater. Power series methods for cellular calculations on atoms, molecules, and solids. In Quantum Science: Methods and Structure: a Tribute to Per-Olov Löwdin [CGLÖ76], pages 57–93. ISBN 0-306-30968-8. LCCN QC174.125 .Q371.

Slater:1977:QTM

[Sla77] John Clarke Slater. Quantum Theory of Matter. Krieger, Huntington, NY, USA, second edition, 1977. ISBN 0-88275-553-6. 763 pp. LCCN QC174.12.

Slater:1978:QTM

[Sla78] John Clarke Slater. Quantum Theory of Molecules and Solids. Volume 4, The Self-consistent Field for Molecules and Solids. Izd-vo “Mir”, Moscow, USSR, 1978. 658 + 4 pp. LCCN 530.1 SL1QT.R.

Slater:1979:CMO

[Sla79] John Clarke Slater. The Calculation of Molecular Orbitals. Wiley, New York, NY, USA, 1979. ISBN 0-471-03181-X. vii + 104 pp. LCCN QD461 .S49 1979. With a foreword by John W. D. Connolly.

Slater:1980:TQD

[Sla80] John C. Slater. Teoria quantistica della materia. (Italian) [Quantum Theory of Matter]. Zanichelli, Bologna, Italy, 1980. ISBN ???? ix + 637 pp. LCCN ????

Slater:1985:WP

[Sla85] John C. Slater. Waves and particles: 1923–1924. In French and Kennedy [FK85], pages 160–162. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Slater:1987:TTK

[Sla87] John C. Slater. Theory of the transition in KH_{2}PO_{4}. Ferroelectrics, 71(1):25–42, 1987. CODEN FEROA8. ISSN 0015-0193 (print),

1563-5112 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00150198708224827. Republication of [Sla41].

Smith:1973:ESR

[SM73] Darwin W. Smith and Walter Bruce McRae, editors. Energy, structure, and reactivity: Proceedings of the 1972 Boulder Summer Research Conference on Theoretical Chemistry. Wiley, New York, NY, USA, 1973. ISBN 0-471-80140-2. LCCN QD462.A1 B68 1972. Proceedings of the Boulder Quantum Theory Symposium and Research Workshop, held at the University of Colorado, Boulder, CO, USA, June 1972.

Smith:1981:JSP

[Smi81] Murphy D. Smith. The John Slater Papers at the American Philosophical Society. Center for History of Physics Newsletter, 13(2): 3–??, ???? 1981. ISSN 0148-5857.

Slater:1969:NON

[SMWW69] J. C. Slater, J. B. Mann, T. M. Wilson, and J. H. Wood. Nonintegral occupation numbers in transition atoms in crystals. Physical Review (2), 184(3):672–694, August 15, 1969. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v184/i3/p672_1.

Sopka:1976:QPA

[Sop76] Katherine Russell Sopka. Quantum physics in America, 1920– 1935. Ph.D. in Education thesis, Graduate School of Arts and Sciences, and the Graduate School of Education, Harvard University, Cambridge, MA, USA, May 1976. various

pp. URL http://hollis.harvard.edu/?itemid=%7Clibrary/m/
aleph%7C000979886;   http://search.proquest.com/docview/

302811033. Reprinted in [Sop80], and revised in [Sop88].

Sopka:1978:BRBa

[Sop78] Katherine R. Sopka. Book review: John C. Slater: Solid-State and Molecular Theory: A Scientific Biography. Isis, 69(2):320–321, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230489.

Sopka:1980:QPA

[Sop80] Katherine Russell Sopka. Quantum Physics in America, 1920– 1935. Three centuries of science in America. Arno Press, New York, NY, USA, 1980. ISBN 0-405-12585-2. xvi + 543 pp. LCCN QC173.98 .S66 1980.

Sopka:1988:QPA

[Sop88] Katherine Russell Sopka. Quantum Physics in America: the years through 1935, volume 10 of The history of modern physics, 1800-1950. Tomash, Los Angeles, CA, USA, 1988. ISBN 0-88318-553-9. xxvi + 376 pp. LCCN QC173.98 .S67 1988. Foreword by Abraham Pais. Introduction by Gerald Holton.

Slater:1936:OAA

[SS36] J. C. Slater and W. Shockley. Optical absorption by the alkali halides. Physical Review (2), 50(8):705–719, October 15, 1936. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v50/i8/p705_1.

Simons:1952:ERB

[SS52] J. C. Simons and J. C. Slater. Electromagnetic resonant behavior of a confocal spheroidal cavity system in the microwave region. Journal of Applied Physics, 23(1):29–30, January 1, 1952. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-

8850. URL http://link.aip.org/link/japiau/v23/i1/p29/s1.

Saffren:1953:APW

[SS53] M. M. Saffren and J. C. Slater. An augmented plane-wave method for the periodic potential problem. II. Physical Review (2), 92(5): 1126–1128, December 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v92/i5/p1126_1.

Slater:1953:TEE

[SSK53] J. C. Slater, H. Statz, and G. F. Koster. A two-electron example of ferromagnetism. Physical Review (2), 91(6):1323–1341, September 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v91/i6/
p1323_1.

Slater:1970:SET

[SW70] John C. Slater and John H. Wood. Statistical exchange and the total energy of a crystal. International Journal of Quantum Chemistry, 5(S4):3–34, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Slater:1971:SET

[SW71] J. C. Slater and J. H. Wood. Statistical exchange and the total energy of a crystal. International Journal of Quantum Chemistry, 4(??):3–??, ?? 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1973:DRF

[SWS^{+}73] John C. Slater, E. Bright Wilson, Darwin W. Smith, Bernard Kirtman, Jerry Goodisman, John W. D. Connolly, and Gayle S. Painter. Discussion remarks following the position paper by J. W. D. Connolly and K. H. Johnson. In Smith and McRae [SM73], pages 106–110. ISBN 0-471-80140-2. LCCN QD462.A1 B68 1972. Proceedings of the Boulder Quantum Theory Symposium and Research Workshop, held at the University of Colorado, Boulder, CO, USA, June 1972.

Slater:1969:CSE

[SWW69] J. C. Slater, Timothy M. Wilson, and J. H. Wood. Comparison of several exchange potentials for electrons in the Cu^{+} ion. Physical Review (2), 179(1):28–38, March 5, 1969. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v179/i1/p28_1.

Slater:19xx:B

[SYLxx] J. C. Slater, H. Yoshizumi, and P.-O. Löwdin. Bibliographies. Uppsala and Florida use only., 19xx.

T:1975:BTM

[T.75] W. J. T. Quantum theory of molecules and solids: Vol. IV, by John C. Slater, McGraw Hill, New York, 1974, pp. xvii + 583, price DM 44.85. Journal of Molecular Structure, 24(2):447– 477, February 1975. CODEN JMOSB4. ISSN 0022-2860 (print),

1872-8014 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022286075870256.

Talman:2004:OID

[Tal04] James D. Talman. Overlap integrals for Dirac–Slater orbitals. International Journal of Quantum Chemistry, 100(2):109–113, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1969:BRP

[Tay69] Philip L. Taylor. Book review: A partisan view: John C. Slater, Quantum Theory of Matter. Physics Today, 22(12):73, 75, December 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v22/i12/p73_s2.

Thomas:1926:MSE

[Tho26] L. H. (Llewellyn Hilleth) Thomas. The motion of the spinning electron. Nature, 117(2945):514, April 10, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Toll:1956:BRB

[Tol56] John S. Toll. Book review: Modern Physics, by John C. Slater. Science, 123(3201):805, May 4, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1750414.

Terrall:1952:PDL

[TS52] J. R. Terrall and John C. Slater. Particle dynamics in the linear accelerator. Journal of Applied Physics, 23(1):66–68, January 1, 1952. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (elec-

tronic), 1520-8850. URL http://link.aip.org/link/japiau/
v23/i1/p66/s1.

Tudela:1993:SRE

[Tud93] David Tudela. Slater’s rules and electron configurations. Journal of Chemical Education, 70(11):956, November 1993. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Uhlenbeck:1925:EHU

[UG25] George E. Uhlenbeck and Samuel Goudsmit. Ersetzung der Hypothese vom unmechanischen Zwang durch eine Forderung bezüglich des inneren Verhaltens jedes einzelnen Elektrons. (German) [Replacement of the hypothesis of unmechanical force with a requirement for the inner behavior of each electron]. Naturwissenschaften, 13(47):953–954, November 1925. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//link.springer.com/article/10.1007/BF01558878;   http:
//www.springerlink.com/content/u20q85248gp8166q/. See [?]

for a computation in 1927 of the radius of a rotating charged particle with the spin of the electron.

Uhlenbeck:1926:SES

[UG26] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. Nature, 117(2938):264–265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). The paper is followed by a half-column supportive letter from Niels Bohr. See [Tho26] for a relativistic correction to the Zeeman splitting of spectral lines.

Utsuno:2013:ECH

[USOA13] Yutaka Utsuno, Noritaka Shimizu, Takaharu Otsuka, and Takashi Abe. Efficient computation of Hamiltonian matrix elements between non-orthogonal Slater determinants. Computer Physics Communications, 184(1):102–108, January 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465512002883.

VanVleck:1964:APC

[Van64] J. H. Van Vleck. American physics comes of age. Physics Today, 17(6):21–??, June 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v17/i6/p21_s1.

VanVleck:1976:OJC

[Van76] John H. Van Vleck. Obituary: John C. Slater. Physics Today, 29 (10):68–69, October 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v29/i10/p68_s1.

VanVleck:1977:RJS

[Van77] John H. Van Vleck. Remarks on John Slater. Slater Memorial Session, American Physical Society, Chicago, 7 February 1977., 1977.

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

Verbeek:1991:GSC

[VV91] Jacob Verbeek and Joop H. Van Lenthe. The generalized Slater– Condon rules. International Journal of Quantum Chemistry, 40(2):

201–210, August 1991. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wagner:1983:CSH

[Wag83] A. J. Wagner. Comparative study of the Hartree–Fock and the Hartree–Fock–Slater method. International Journal of Quantum Chemistry, 23(1):235–248, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wasserman:1981:BKS

[Was81] Neil H. Wasserman. The Bohr–Kramers–Slater Paper and the Development of the Quantum Theory of Radiation in the Work of Niels Bohr. Ph.D dissertation, Harvard University, Cambridge, MA, USA, 1981.

Weiner:1970:OHI

[Wei70a] Charles Weiner. Oral history interview with John Clarke Slater, 1970 February 23 and August 7. Web transcript., 1970. URL http:

//www.aip.org/history/ohilist/4893_1.html.

Weiner:1970:PGD

[Wei70b] Charles Weiner. Physics in the Great Depression. Physics Today, 23(10):31–38, October 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v23/i10/p31_s1. Reprinted in [WP85,

pages 115–121].

Weisstein:2007:SJC

[Wei07] Eric W. Weisstein. Slater, John Clarke (1900–1976). Wolfram Research ScienceWorld online., 2007. URL http://scienceworld.

wolfram.com/biography/Slater.html.

Weller:1967:FPS

[Wel67] Edward F. Weller, editor. Ferroelectricity: proceedings of the Symposium on Ferroelectricity, General Motors Research Laboratories, Warren, Michigan, 1966. Elsevier, Amsterdam, The Netherlands, 1967. LCCN QC595 S98.

Waldron:2001:SPB

[WFS^{+}01] Kimberley A. Waldron, Erin M. Fehringer, Amy E. Streeb, Jennifer E. Trosky, and Joshua J. Pearson. Screening percentages based on Slater effective nuclear charge as a versatile tool for teaching periodic trends. Journal of Chemical Education, 78(5):635, May 2001. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Wilson:1976:OSJ

[Wil76] H. W. Wilson. Obituary: Slater, John Clarke, 1900–1976 (American physicist). New York Times, page 32, July 27, 1976. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Witmer:1952:BRB

[Wit52] Enos E. Witmer. Book review: Quantum Theory of Matter, by John C. Slater. American Journal of Physics, 20(4):251, April 1952. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v20/
i4/p251_s1.

Wood:1962:EBI

[Woo62] J. H. Wood. Energy bands in iron via the Augmented Plane Wave method. Physical Review (2), 126(2):517–527, April 1962. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://link.aps.org/doi/10.1103/PhysRev.

126.517. See [Sla37e] for the original APW method.

Wood:1957:WFE

[WP57] J. H. Wood and G. W. Pratt. Wave functions and energy levels for Fe as found by the Unrestricted Hartree–Fock method. Physical Review (2), 107(4):995–1001, August 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.107.995.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of Physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Wilson:1970:SSE

[WWS70] Timothy M. Wilson, J. H. Wood, and J. C. Slater. Studies of the statistical exchange approximation in first-transition-row atoms and ions: Mn^{+2} ion. Physical Review A (Atomic, Molecular, and Optical Physics), 2(3):620–634, September 1, 1970. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

Zhang:2011:CGS

[ZD11] Yong Zhang and Xuanchun Dong. On the computation of ground state and dynamics of Schrödinger–Poisson–Slater system. Journal of Computational Physics, 230(7):2660–2676, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000076.